### v1

# Towards Automatic Evaluation and High-Quality Pseudo-Parallel Dataset Construction for Audio Editing: A Human-in-the-Loop Method

Abstract
Audio editing aims to manipulate audio content based on textual descriptions, supporting tasks such as adding, removing, or replacing audio events. Despite recent progress, the lack of high-quality benchmark datasets and comprehensive evaluation metrics remains a major challenge for both assessing audio editing quality and improving the task itself. In this work, we propose a novel approach for audio editing task by incorporating expert knowledge into both the evaluation and dataset construction processes: 1) First, we establish AuditScore, the first comprehensive dataset for subjective evaluation of audio editing, consisting of over 6,300 edited samples generated from 7 representative audio editing frameworks and 23 system configurations. Each sample is annotated by professional raters on three key aspects of audio editing quality: overall Quality, Relevance to editing intent, and Faithfulness to original features. 2) Based on this dataset, we train AuditEval, the first model designed for automatic MOS-style scoring tailored to audio editing tasks. AuditEval addresses the critical lack of objective evaluation metrics and the prohibitive cost of subjective assessment in this field. 3) We further leverage AuditEval to evaluate and filter a large amount of synthetically mixed editing pairs, constructing a high-quality pseudo-parallel dataset by selecting the most plausible samples. Objective experiments validate the effectiveness of our expert-informed filtering strategy in yielding higher-quality data, while also revealing the limitations of relying solely on objective metrics. The dataset, codes and tools can be found at: https://github.com/NKU-HLT/AuditEval.
Introduction
Benefiting from the construction of Text-to-Audio (TTA) and Audio Caption (AC) datasets(Kim et al. 2019; Gemmeke et al. 2017; Drossos, Lipping, and Virtanen 2020; Mei et al. 2024; Piczak 2015; tza 2002), audio generation has made significant strides in recent years. Current TTA models—spanning diffusion-based frameworks(Huang et al. 2023; Liu et al. 2023, 2024b; Ghosal et al. 2023; Chen et al. 2024; Majumder et al. 2024; Hung et al. 2024; Xue et al. 2024) and autoregressive Transformer-based systems(Agostinelli et al. 2023; Kreuk et al. ; Copet et al. 2023)—are capable of producing high-quality, diverse, and text-aligned audio outputs based on user prompts. However, beyond generation, audio editing remains equally essential and presents more formidable challenges. Unlike generation, which involves creating audio from scratch, audio editing requires the system to modify existing audio content according to user-provided textual instructions or target descriptions. Unlike generation, which creates audio from scratch, editing modifies existing content based on textual instructions or target descriptions. It includes basic event manipulations (e.g., adding, removing, replacing) and more advanced tasks like super-resolution and style transfer. Audio editing poses three major challenges: maintaining audio quality, ensuring precise control aligned with user intent, and preserving the unedited portions of the original audio, rather than generating an entirely new sample.
However, the development of audio editing has been significantly hindered by the lack of high-quality paired datasets—an issue also observed in the vision domain. Constructing supervised audio editing datasets is particularly difficult: unlike text-to-audio data, naturally occurring audio pairs that reflect specific editing operations are virtually nonexistent. Manual collection is often prohibitively expensive or infeasible—especially for scenarios like replacing gunshots in a crowd with thunder while preserving alignment. As a result, most existing methods adopt training-free approaches based on generative diffusion models(Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) and inversion techniques(Song, Meng, and Ermon 2020; Mokady et al. 2023; Wu and De la Torre 2023; Huberman-Spiegelglas, Kulikov, and Michaeli 2024). These methods repurpose the generation and cross-modal alignment capabilities of pretrained TTA models, steering generation toward desired edits by adjusting prompts or conditioning signals(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a; Tsai et al. 2024). However, they heavily depend on the underlying TTA model and often lack robustness due to the absence of task-specific supervision.
As an alternative to training-free approaches, supervised frameworks have also been explored by training models on explicitly constructed editing datasets. Audit(Wang et al. 2023b) represents the first supervised audio editing framework, introducing a triplet-based dataset and training a dedicated latent diffusion model (LDM). To date, it remains the only task-specific paired dataset for audio editing. This dataset is created through audio mixing and combinatorial pairing strategies, leveraging the compositional nature of audio—unlike images, audio signals can be naturally combined to simulate coherent edits. A similar construction paradigm is adopted in InstructME(Han et al. 2023), which applies mixing-based strategies to generate supervision signals for music editing tasks. Meanwhile, MusicGen(Copet et al. 2023) and Instruct-MusicGen(Zhang et al. 2024a) explore autoregressive approaches based on discrete audio-text tokens, supporting tasks such as continuation, transformation, and editing. WavCraft(lia 2024) further investigates instruction-based editing by employing large language models (LLMs) to interpret textual commands and orchestrate external audio and speech models for modification.
Despite recent progress, there is still no unified standard for evaluating audio editing tasks. The absence of a widely accepted, open-source benchmark specifically designed for editing has led to fragmented evaluation protocols and a lack of consistent objective metrics. Currently, the most commonly used metric is the CLAP Score, which measures similarity between the edited audio and the target prompt in the embedding space, leveraging the CLAP model(Elizalde et al. 2023) for cross-modal comparison without requiring ground-truth audio. To assess preservation of unedited regions, some works(Paissan et al. 2023; Zhang et al. 2024b; Jia et al. 2025; Liu et al. 2024a) use the original audio as a reference and compute similarity between the original and edited versions. Others(Xu et al. 2024; Zhang et al. 2024b) construct pseudo-paired datasets via audio mixing—following a strategy similar to Audit and InstructME—and evaluate outputs by comparing them to mixed signals using conventional TTA metrics such as Fréchet Distance (FD), Fréchet Audio Distance (FAD)(Roblek et al. 2019), and KL divergence. However, such evaluation paradigms—relying on synthetic mixtures as ground truth—often fall short. Since high-quality audio edits may vary significantly and are not strictly aligned with a single reference, these metrics can fail to reliably capture true editing quality or model controllability.
As a result, subjective human evaluation remains the most reliable method for assessing audio editing quality. Current subjective evaluation protocols typically consider three aspects: overall audio Quality(Zhang et al. 2024b; Jia et al. 2025), Relevance to the editing intent(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a), and Faithfulness to the original (unedited) content(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a). These criteria provide an intuitive and comprehensive assessment of audio editing performance. Nevertheless, compared to pure generation tasks such as Text-to-Audio (TTA), Text-to-Speech (TTS), or Text-to-Music (TTM), audio editing is inherently more complex, requiring deeper understanding and recognition of nuanced content. This results in increased costs for annotator training, longer evaluation durations, and greater human labor requirements in practice. In light of these challenges, developing automatic evaluation methods and building a dedicated evaluation system for audio editing is both meaningful and valuable for advancing research in this field.
In this paper, we aim to bridge the evaluation gap in audio editing by introducing AuditEval, a scalable automatic MOS-style assessment framework, and subsequently leveraging it to build a high-quality pseudo-parallel dataset. Our contributions are threefold:
  • Subjective Audio Editing Evaluation Dataset Construction. We construct AuditScore, the first comprehensive benchmark for subjective audio editing evaluation, consisting of over 6,300 edited samples from 7+ representative frameworks and 23+ system variants. Each sample is annotated by professional raters on three key aspects: Quality, Relevance, and Faithfulness.
  • Automatic MOS Prediction. Based on AuditScore, we train AuditEval, the first model tailored for automatic MOS-style scoring in audio editing. It effectively mitigates the lack of objective metrics and the high cost of human evaluation.
  • High-Quality Pseudo-Parallel Audio Editing Dataset Construction. We further utilize AuditEval to assess and filter a large volume of synthetically mixed editing pairs, selecting only the most plausible samples to construct a high-quality pseudo-parallel dataset. Objective experiments validate the effectiveness of this expert-informed filtering strategy in producing superior-quality data, while also highlighting the shortcomings of relying exclusively on objective metrics.
Related Work
Training-free Audio Editing Method
Training-free audio editing methods exploit pre-trained TTA diffusion models to perform editing without the need for task-specific fine-tuning. These approaches typically begin by inverting the input audio into the noise space of the diffusion model—that is, mapping from $x_{0}$ (the observed audio) to $x_{T}$ (a noisy latent representation). The resulting latent $x_{T}$ preserves the semantic and structural features of the original audio while providing a flexible representation for guided generation. Editing is then achieved by denoising $x_{T}$ into a modified output $\hat{x}_{0}$, conditioned on user instructions or target descriptions. Depending on the specific inversion strategy used, training-free audio editing methods can be broadly categorized into three types: SDEdit-based, DDIM-based, and DDPM-based approaches.
SDEdit-based methods perform audio editing by first applying a forward stochastic differential equation (SDE) process to produce a noised version of the original audio, and then conducting denoising under a modified prompt using the reverse SDE process (Eq. 1). In this formulation, $\alpha(t)$ and $\sigma(t)$ are time-dependent functions that govern the noise injection level, and $\boldsymbol{s}_{\theta}$ represents the pretrained score model. The inversion timestep $t\in[0,T]$ plays a crucial role: larger $t$ values (closer to $T$) lead to more realistic yet less faithful edits, while smaller $t$ values better preserve the original content but reduce the editing strength.
$\mathbf{x_{t}}={}$ $\mathbf{x_{t+\Delta t}}+\left(\textstyle\sigma^{2}(t)\!-\!\sigma^{2}(t+\Delta t)\right)\cdot$ (1)
$\boldsymbol{s}_{\theta}(\mathbf{x_{t}},t)+\sqrt{\textstyle\sigma^{2}(t)\!-\!\sigma^{2}(t+\Delta t)}\,\mathbf{z}$
DDIM-based approaches perform deterministic inversion of the diffusion process by exploiting the approximate reversibility of the underlying ODE with sufficiently small step sizes (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Mokady et al. 2023). This allows for editing by first reversing the generation trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}T$ (Eq. 2, where $\alpha$ is a hyperparameter and $\varepsilon{\theta}$ denotes the pretrained latent diffusion model), and then regenerating the audio using a modified prompt (Eq. 3), thereby achieving semantically guided edits. Building upon DDIM inversion, several studies (Xu et al. 2024; Zhang et al. 2024b; Jia et al. 2025; Liu et al. 2024a) incorporate enhancements such as attention map manipulation and null-text optimization, demonstrating effective results in controllable audio editing.
$x_{t+1}=\textstyle\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}x_{t}+\left(\sqrt{\frac{1-\alpha_{t+1}}{\alpha_{t+1}}}\!-\!\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}\right)\cdot\varepsilon_{\theta}(x_{t},t,\mathcal{P}_{\boldsymbol{ori}})$ (2)
$x_{t-1}=\textstyle\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}x_{t}+\left(\sqrt{\frac{1-\alpha_{t-1}}{\alpha_{t-1}}}\!-\!\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}\right)\cdot\varepsilon_{\theta}(x_{t},t,\mathcal{P}_{\boldsymbol{tar}})$ (3)
DDPM-based approaches include CycleDiffusion(Wu and De la Torre 2023) and Edit-Friendly DDPM(Huberman-Spiegelglas, Kulikov, and Michaeli 2024). This kind of method leverages the stochastic nature of the DDPM sampling process—specifically, the stochastic noise injected at each diffusion step—to preserve the characteristics of the original input. By adding noise to the original input and storing intermediate latent states at each step, the model can compute the corresponding stochastic noise components (Eq.4, $\mu_{t}\left(\mathrm{x}_{t}\right)$ is the calculated expectation of ${x}_{t}$, $\sigma_{t}$ is the hyperparameter), which are then reused during the denoising process under the modified prompt to retain original content. Edit-Friendly DDPM further improves by altering the noise injection mechanism so that the noise added at each timestep is independent, thereby enhancing editing controllability and performance(Huberman-Spiegelglas, Kulikov, and Michaeli 2024). ZETA Editing(Manor and Michaeli 2024) is the first known method to apply DDPM inversion to audio editing, demonstrating superior results compared to SDEdit- and DDIM-inversion-based methods.
$\mathbf{z}_{t}=\left(\mathrm{x}_{t-1}-\mu_{t}\left(\mathrm{x}_{t}\right)\right)/\sigma_{t},\quad t=T,\ldots,1$ (4)
Supervised Audio Editing Models
In contrast to training-free methods, supervised audio editing approaches rely on explicitly constructed editing datasets and can be broadly divided into non-autoregressive and autoregressive paradigms. Models such as Audit and InstructME employ latent diffusion models conditioned on audio-text multimodal inputs. They utilize specialized conditioning mechanisms and position-aware loss functions to enable end-to-end training on pseudo-labeled editing pairs or instrument transfer datasets, achieving effective editing performance (Wang et al. 2023b; Han et al. 2023). On the other hand, autoregressive approaches generate edited audio by conditioning on discrete audio-text representations of the original content and edit instructions as context. Representative models like MusicGen and Instruction-MusicGen adopt transformer-based decoders to sequentially synthesize edited audio (Copet et al. 2023; Zhang et al. 2024a). While these models benefit from strong language-audio alignment and high generation quality, they often struggle with temporal inconsistencies between the original and edited audio.
Automatic Evaluation for Speech, Music and Audio
Automatic subjective evaluation plays a vital role in assessing AI-generated outputs by simulating human perception in a scalable and repeatable manner. Compared to purely objective metrics, it provides a more intuitive and comprehensive measure of perceptual quality. A well-designed evaluator not only aligns with human preferences but also guides the development of future generative models. In the speech domain, automatic Mean Opinion Score (MOS) prediction has been widely adopted to evaluate tasks such as Text-to-Speech (Cooper et al. 2022; Wang et al. 2023a, 2024), Text-to-Voice (Tang et al. 2024), and speech enhancement(Reddy, Gopal, and Cutler 2021; Mittag et al. 2021; Chen et al. 2022). These models have shown promising results and effectively complement traditional objective metrics.
In contrast, automatic evaluation for audio and music remains more challenging due to the greater content diversity—from natural sound events to complex musical structures—and the inherently looser alignment between audio and textual descriptions compared to the tighter correspondence in speech-transcription pairs. To address these challenges, MusicEval(Liu et al. 2025) is the first known framework to provide automatic scoring for music generation. It constructs a high-quality evaluation dataset and trains scoring models to assess both overall quality and text relevance, achieving strong correlation with human judgments. Meanwhile, Meta’s Audiobox Aesthetics(Tjandra et al. 2025) proposes a unified framework for speech, music, and audio evaluation, defining four core aesthetic dimensions that collectively provide a comprehensive and structured standard for automatic audio assessment from an aesthetic perspective.
Data Collection
Basic Information of AuditScore
Table 1: Subjective evaluation result. Detailed distribution of difference editing types is presented in AppendixH I.
Editing
Model or Method
Generative
TTA Backbone
Timesteps
System
ID
Mean Scores
Quality
Relevance
Faithfulness
SDEdit Tango2 75 sys1 3.621 2.896 3.699
125 sys2 3.622 3.420 3.375
175 sys3 3.548 3.503 2.893
Auffusion 75 sys4 3.602 3.268 3.426
125 sys5 3.487 3.461 2.885
175 sys6 3.492 3.531 2.742
DDIM Inversion Tango2 200 sys7 3.420 2.791 3.489
Auffusion 200 sys8 3.867 3.447 3.441
Edit-friendly
DDPM Inversion
Tango2 75 sys9 3.658 3.330 3.641
125 sys10 3.684 3.354 3.479
175 sys11 3.655 3.593 3.408
Auffusion 75 sys12 3.639 3.029 3.483
125 sys13 3.687 3.431 3.417
175 sys14 3.669 3.432 3.167
CycleDiffusion
DDPM Inversion
Tango2 75 sys15 3.592 2.848 3.595
125 sys16 3.445 3.184 3.313
175 sys17 3.293 3.146 3.054
Auffusion 75 sys18 3.463 2.922 3.423
125 sys19 3.360 3.103 3.047
175 sys20 3.262 3.159 2.893
AP-adapter Audioldm2 - sys21 3.353 2.857 2.653
Musicgen - - sys22 3.208 2.968 2.693
Human Edit - - sys23 3.883 3.878 3.988
AuditScore is a human-annotated dataset comprising 6,360 pairs of original and edited audio samples, totaling 35.3 hours in duration. These samples are generated using 23 distinct system configurations across seven representative audio editing approaches: SDEdit (Meng et al. ), DDIM (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020), Edit-Friendly DDPM (Huberman-Spiegelglas, Kulikov, and Michaeli 2024; Manor and Michaeli 2024), CycleDiffusion (Wu and De la Torre 2023), AP-adapter (Tsai et al. 2024), MusicGen (Copet et al. 2023), and Human Edit(manual editing). For diffusion-based methods (SDEdit, DDIM, Edit-Friendly DDPM, and CycleDiffusion), we adopt two state-of-the-art generative backbones—Tango2 (Majumder et al. 2024) and Auffusion (Xue et al. 2024)—to synthesize the edited audio. To further promote diversity and enhance model generalization, each configuration is instantiated with three inversion timesteps (75, 125, and 175 out of 200), allowing finer control over the degree of editing. This setup is specifically applied to SDEdit, Edit-Friendly DDPM, and CycleDiffusion to explore a wider range of editing intensities and semantic variations.
To construct a representative and diverse dataset, we began by selecting original audio samples and corresponding captions from the AudioCaps dataset (Kim et al. 2019). Captions were clustered into 100 semantic groups using GTE-Large text embeddings (Li et al. 2023), from which 2–3 audio samples were manually selected per cluster. For each selected audio, we crafted three types of editing instructions—Addition, Deletion, and Replacement—yielding a shared pool of 240 instructions across all systems. Additionally, several representative systems were assigned 30 unique instructions to further increase variation in editing scenarios. In total, AuditScore covers over 200 distinct audio event types, with editing operations introducing more than 150 novel events. The rich variety of methods, model architectures, editing parameter settings, and content diversity enables AuditScore to capture a wide spectrum of editing behaviors, audio qualities, and modeling challenges—making it a comprehensive and robust benchmark for training and evaluating audio editing evaluation models.
To assess the edited audio samples generated by these 23 systems, we conducted a large-scale subjective listening study involving five domain experts with extensive experience in audio dataset curation and perceptual evaluation. Each expert independently rated the editing audio pairs along three key perceptual dimensions: Quality, Relevance, and Faithfulness. The final rating for each system was computed as the average of expert scores across all three dimensions. A comprehensive summary of system IDs and their evaluation results is presented in Table 1.
(a) Quality (Distribution)
(b) Relevance (Distribution)
(c) Faithfulness (Distribution)
(d) Quality (Heatmap)
(e) Relevance (Heatmap)
(f) Faithfulness (Heatmap)
Figure 1: Comparison of quality, relevance, and faithfulness across all systems and annotation types. Top: score distribution across samples. Bottom: system-level heatmaps from AuditScore.
Scoring Protocol and Analysis
To comprehensively evaluate audio editing performance, we adopt a three-dimensional scoring protocol encompassing Quality, Relevance, and Faithfulness. These dimensions are informed by established perceptual evaluation methodologies—such as MOS (Mean Opinion Score), CMOS (Comparison MOS)(Loizou 2011), and semantic alignment assessments—while tailored specifically for the nuances of audio editing. Detailed annotation guidelines were developed to ensure consistency across expert raters while preserving flexibility for nuanced perceptual judgments. The scoring scales are defined as follows(Additional scoring rubrics and details can be found in AppendixA B):
  • Quality — Assesses the overall semantic correctness and acoustic integrity of the edited audio: (1) About the Same / Even Better, (2) Slightly Worse, (3) Noticeably Worse, (4) Much Worse, (5) Severely Degraded.
  • Relevance — Measures how well the edited audio semantically matches the editing instruction and target description: (1) Excellent Match, (2) Good Match, (3) Partial Match, (4) Poor Match, (5) Mismatch.
  • Faithfulness — Evaluates the preservation of unedited content from the original audio in terms of rhythm, timing, loudness, and overall style: (1) Perfectly Faithful, (2) Mostly Faithful, (3) Partially Faithful, (4) Minimally Faithful, (5) Not Faithful.
Figure 1a–c show the distribution of scores across the three dimensions. The scores for Quality, Relevance, and Faithfulness each approximately follow a normal distribution, indicating that our scoring criteria provide a balanced and well-calibrated assessment scale. Notably, quality scores are slightly skewed toward higher values, reflecting that most editing systems tend to preserve audio quality easily, while still facing challenges in precisely executing editing instructions and maintaining content fidelity.
Figure 1b–d further visualize performance across different systems. Notably, Human Editing consistently achieves the highest ratings in all dimensions, establishing an upper bound for automated models. Interestingly, some systems (e.g., sys3, sys5, and sys6) achieve high relevance scores, indicating strong alignment with the editing instructions, yet score lower in faithfulness—highlighting a common trade-off in current models between effective editing and content preservation. This analysis underscores the remaining gap between human and machine editing and suggests a clear direction for future model improvement.
Figure 2: The overall architecture of AuditEval.
Table 2: Performance of the proposed model on the AuditScore dataset at system and utterance levels across three dimensions.
Dimension System-level Utterance-level
MSE↓ LCC↑ SRCC↑ KATU↑ MSE↓ LCC↑ SRCC↑ KATU↑
Quality Score 0.0595 0.6101 0.7004 0.5516 0.1474 0.3318 0.3306 0.2462
Relevance Score 0.0962 0.6196 0.6472 0.4783 0.2963 0.4105 0.3991 0.2886
Faithfulness Score 0.0726 0.8460 0.8809 0.7024 0.2217 0.5327 0.5171 0.3791
Table 3: Comparison of pseudo-parallel data quality before and after filtering. Objective metrics include CLAPt (cross-modal similarity between edited audio and target text), CLAPa(similarity between edited audio and target audio), IS, FAD, FD, and KL divergence (all measuring distributional similarity between edited and original audio). Subjective metrics are predicted human preference scores from AuditEval, used to compare perceived quality and reveal discrepancies with objective measures.
Dataset Objective Subjective (Automatic)
CLAPt(%)↑ CLAPa(%)↑ IS↑ FAD↓ FD↓ KL↓ Quality Relevance Faithfullness
Random1 48.70 82.73 12.23 0.2859 3.207 0.6913 3.365 3.281 3.114
Random2 48.47 82.69 12.23 0.2816 3.197 0.6892 3.365 3.283 3.114
Random3 48.76 82.73 12.20 0.2830 3.205 0.6867 3.366 3.281 3.114
Filtered 49.43 82.76 12.61 0.4587 4.356 0.6783 3.626 3.694 3.493
Method
AuditEval: An Automatic Evaluation Model
Based on our newly curated AuditScore dataset, we develop AuditEval—an automatic MOS-style evaluation model for audio editing built upon the MusicEval framework. AuditEval processes two groups of inputs: the original audio paired with its edited version, and the corresponding textual descriptions of the audio before and after editing, respectively. It outputs three scalar scores reflecting the key evaluation dimensions: Quality, Relevance, and Faithfulness. To account for the different learning complexities across these dimensions, AuditEval employs three independently trained decoding heads, each specialized for predicting one specific score. This simple yet effective design represents a minimal adaptation of the MusicEval framework (see Figure 2), with modest architectural changes and some optimization efforts, while still achieving balanced and reliable predictions across all evaluation aspects.
Pseudo-parallel Data Filtering
Due to the lack of well-established supervised datasets for audio editing, we construct a large-scale (500,000 pairs, 2700h) Pseudo-parallel Audio Editing Dataset following the methodologies of AUDIT (Wang et al. 2023b) and InstructME (Han et al. 2023), and further enhance it using our automatic scoring model, AuditEval. The dataset consists of synthetic audio tuples (A, B, C, A+B, A+C), where Audio A is sampled from clustered AudioCaps (Kim et al. 2019) descriptions, and B/C are single-event clips paired with semantically similar prompts. These are reorganized into instruction-based triplets for Addition, Deletion, and Replacement tasks. A rule-based strategy is used to mix the audio and construct corresponding textual instructions, aiming to maintain acoustic coherence and semantic plausibility. Details are provided in AppendixB C.
However, this kind of rule-based concatenation introduces several limitations: (1) the quality and prompt relevance of individual audio elements are inconsistent; (2) the insertion positions in the background audio are often suboptimal—even when attempting to insert at quieter regions, the resulting mixtures can sound unnatural and make it difficult to distinguish the editing change; (3) the randomly selected A/B/C combinations can be semantically implausible—for example, inserting the sound of paper rustling into a battlefield recording, or replacing male speech with a toilet flushing sound, etc. To address these issues, we apply AuditEval to automatically evaluate and filter synthesized audio editing pairs based on scoring criteria informed by AuditScore. Because AuditScore covers almost all mainstream editing operations including mixing, this filtering is expected to improve data quality and semantic plausibility, advancing toward a practical and scalable supervised dataset (see AppendixC D).
Experiments
Dataset and Metrics
For the automatic evaluation of audio editing, we employ our proposed AuditEval model and train it on the curated AuditScore dataset using a fixed 9:1 train-test split. The dataset will be publicly released to facilitate reproducibility. We report the Mean Squared Error (MSE) at both the utterance and system levels to quantify prediction accuracy. To evaluate consistency with human judgments, we additionally report the Linear Correlation Coefficient (LCC), Spearman’s Rank Correlation Coefficient (SRCC), and Kendall’s Tau Correlation Coefficient (KTAU)(Cooper et al. 2022), which capture linear and rank-based relationships between predicted scores and ground-truth annotations. For the task of selecting high-quality pseudo-parallel audio editing data, we conduct exploratory experiments with objective evaluations to compare the audio quality before and after filtering. Referring to typical training data scales, we retain the top 15% of data for comparison. Specifically, we adopt FAD, FD, and KL, following the methodology of (Liu et al. 2023), to assess the distributional similarity between audio sets. Additionally, we use the CLAP Score to evaluate cross-modal alignment between audio and corresponding descriptions.
(a) Quality
(b) Relevance
(c) Faithfulness
Figure 3: System-level comparison between AuditEval predictions and human ratings across three evaluation aspects.
Results and Analysis
We first evaluate the effectiveness of our automatic evaluation model, AuditEval, which is trained on the curated AuditScore dataset. Table 2 presents the evaluation results at both the system and utterance levels.
At the system level, the mean squared error (MSE) between AuditEval predictions and human ratings remains impressively low—below 0.1 on a 1–5 scale—across all evaluation dimensions. Meanwhile, the linear correlation coefficient (LCC) exceeds 0.6 in every case. Notably, Faithfulness achieves an LCC of 0.84, indicating strong agreement with human judgments. As illustrated in Figure 3, AuditEval effectively ranks different models according to their performance, particularly in terms of executing editing intentions (Relevance) and preserving unedited content (Faithfulness). These results demonstrate that AuditEval successfully fulfills its goal of bridging the evaluation gap in audio editing by providing a scalable and repeatable MOS-style assessment tool to support model comparison and benchmarking.
At the utterance level, the mean squared error (MSE) between AuditEval predictions and human ratings remains below 0.3 (on a 1–5 scale) across all dimensions, with the linear correlation coefficient (LCC) exceeding 0.3 in each case, indicating positive correlation (see Figure 3). Compared to system-level evaluation, utterance-level prediction is inherently more challenging due to increased variability. Nevertheless, AuditEval achieves relatively low MSE values and maintains a reasonable correlation with human judgments, demonstrating its ability to approximate human perception at finer granularity. Importantly, the closeness of AuditEval’s predictions to human ratings in absolute terms lays a solid foundation for downstream applications such as automatic utterance-level filtering or selection.
Among the three dimensions, Faithfulness consistently achieves the highest performance, both at the system and utterance levels. This indicates that the model is relatively effective in evaluating whether the unedited content is preserved after editing. In contrast, Relevance and Quality are more challenging to predict accurately, as they require a deeper joint understanding of both the original and edited audio-text pairs. Unlike standard audio generation evaluation—which typically involves assessing a single audio clip against its corresponding text—audio editing evaluation introduces an additional layer of semantic and acoustic reasoning across paired inputs.
To further assess the effectiveness of our proposed filtering strategy for the Pseudo-parallel Audio Editing Dataset, we compare the quality of three randomly sampled subsets (Random1-3) against a curated subset (Filtered) using standard objective metrics. As shown in Table 3, the Filtered subset outperforms three random baselines across most of the metrics, achieving the highest CLAP-based semantic alignment (CLAP_t: 49.43%, CLAP_a: 82.76%), as well as superior IS (12.61) and KL (0.6783) scores.
(a) Quality
(b) Relevance
(c) Faithfulness
Figure 4: Utterance-level fit between AuditEval predictions and human ratings cross three evaluation aspects.
These results demonstrate that the proposed filtering process significantly improves semantic consistency and editing quality in the pseudo-parallel data.
Nevertheless, certain objective quality metrics, such as FAD and FD, do not fully capture the improvements brought by our filtering strategy. Interestingly, the Filtered subset exhibits noticeably worse FAD and FD scores compared to the random subsets. This discrepancy underscores a fundamental divergence in the focus of subjective and objective evaluations: while human judgments are more attuned to high-level semantic attributes—such as relevance, coherence, and editing accuracy—objective metrics like FAD and FD measure more low-level acoustic properties and may even surpass human perception in such fine-grained distinctions. This distinction explains why semantically grounded metrics (e.g., CLAP, IS) align more closely with human evaluations, whereas acoustic metrics tend to diverge. Therefore, subjective evaluation remains a vital and complementary tool for assessing editing quality, particularly in semantically rich tasks.
Conclusion
In this paper, we address the longstanding challenge of evaluating audio editing systems by introducing AuditEval, a scalable framework for automatic MOS-style assessment. Through the construction of the AuditScore benchmark and the development of the AuditEval model, we enable accurate and efficient quality estimation across multiple dimensions. Furthermore, we demonstrate the practical utility of AuditEval by applying it to curate a high-quality pseudo-parallel dataset, which effectively leads to tangible improvements in data quality. Overall, our work lays a solid foundation for future research in both evaluation and data construction for audio editing.
Appendix A Appendix
Appendix B A. Details of Scoring Criteria Specification
Here, we present details about Scoring Criteria Specification. As described in the main text, we evaluate audio editing quality across three dimensions: Quality, Relevance, and Faithfulness. we establish detailed scoring criteria to ensure consistent and meaningful annotations. These guidelines integrate established evaluation practices including MOS for perceived quality, CMOS for relative assessment, and text-audio alignment for semantic consistency. The detailed specifications for these three evaluation dimensions are presented below. Each dimension is scored on a 5-point scale (1-5), where higher values indicate better performance.
Quality score reflects how well the edited audio aligns with the editing instruction and target description, in terms of semantic correctness and execution:
$\bullet$ 5 - About the Same/Even Better: The edited audio is on par with or even better than the original in terms of intelligibility, fluency, and background noise.
$\bullet$ 4 - Slightly Worse: At most one aspect (e.g., mild mechanical artifact or minor increase in noise) is slightly degraded, but overall listening quality remains unaffected.
$\bullet$ 3 - Noticeably Worse: At least two aspects show moderate degradation, or one aspect is clearly degraded. Artifacts such as robotic tone, discontinuities, or increased noise may affect understanding in some segments but remain tolerable.
$\bullet$ 2 - Much Worse: Significant degradation in at least two aspects, with noticeable artifacts severely affecting naturalness and intelligibility.
$\bullet$ 1 - Badly Damaged: Severe issues across all aspects, such as overwhelming noise, harsh distortions, or unintelligible content, rendering the audio difficult to comprehend.
Relevance score reflects how well the edited audio aligns with the editing instruction and target description, in terms of semantic correctness and execution:
$\bullet$ 5 - Excellent Match: The edit perfectly matches the described intent. For Addition, the added event feels natural and contextually appropriate. For Deletion, the targeted event is successfully removed without leaving any noticeable residue. For Replacement, the substituted content is fluent and coherent, maintaining the overall auditory experience.
$\bullet$ 4 - Good Match: Overall alignment is correct, but there may be minor imperfections, such as slightly abrupt additions or faint remnants after deletion.
$\bullet$ 3 - Partial Match: The general category or type of edit is correct, but details may be awkward or somewhat misaligned with the description.
$\bullet$ 2 - Poor Match: The edit loosely corresponds to the intent, but contains semantic inaccuracies or unnatural execution that hampers comprehension.
$\bullet$ 1 - Mismatch: The edited content fails to reflect the described intent, with missing, incorrect, or misleading modifications.
Faithfulness criterion evaluates how well the unedited portions of the audio are preserved in terms of rhythm, sequence, loudness, and overall style:
$\bullet$ 5 - Perfectly Faithful: Unedited segments are completely preserved. Prosody, order, and audio style remain fully consistent with the original.
$\bullet$ 4 - Mostly Faithful: Slight variation in loudness or subtle stylistic shifts may occur, but prosody and content are largely preserved.
$\bullet$ 3 - Partially Faithful: Noticeable changes in rhythm or sequence are present, though the overall style remains similar. Minor alterations in the unedited content may be perceptible.
$\bullet$ 2 - Low Faithfulness: Major changes in prosody, ordering, or style, with only a few elements traceable to the original.
$\bullet$ 1 - Not Faithful: The edited audio bears little or no resemblance to the original; unedited content is altered significantly or missing altogether.
Appendix C B. Details of Pseudo-Parallel Audio Editing Dataset Construction.
Inspired by AUDIT and InstructME, we also adopt an audio mixing strategy to construct our Pseudo-Parallel Audio Editing Dataset. The goal of the dataset construction process is to generate a set of audio pairs comprising (A, B, C, A + B, A + C), which serves as the foundation for subsequent audio editing tasks. For audio types A, we collected over 45,000 audio prompts from the AudioCaps dataset (Kim et al. 2019). For audio types B and C, we select prompts corresponding to common single-audio events, covering eight broad categories—NaturalSounds, Transportation, IndoorActivities, OutdoorActivities, Animals, Music, Speech and HumanSounds—spanning over 100 distinct sound types. By pairing these prompts in all possible two-way combinations, we obtained 4950 pairs of audio prompts. Then, for each audio prompt pair (B, C), we randomly select 20 audio prompts from the 45,000+ A audio prompts to form combinations, denoted as (A, B, C).
We utilize TangoFlux(Hung et al. 2024) to generate audio based on the (A, B, C) prompts, using a step of 50 and a guide scale of 4.5. The generation rules are as follows:
  • Audio A: The generated audio has a fixed length of 10 seconds.
  • Audio B: The duration of the generated audio is randomized, with a duration ranging from 3 to 6 seconds.
  • Audio C: The duration of the generated audio matches that of the audio B.
Then we present the method used to mix the audio A, B, and C to obtain the audio A+B and A+C. Taking the generation of A+B as an example, let the duration of audio B be denoted as $t_{B}$. We apply a sliding window approach to identify the segment in audio A with the lowest energy and a length equal to $t_{B}$, which is then selected as the insertion interval for audio B. To reduce the search time, the stride of the sliding window was set to 0.1 seconds. To ensure consistency of volume and alignment of the dynamic range between the two audio segments, the amplitude of audio B is adjusted. Specifically, peak amplitude normalization is applied to audio B so that its maximum amplitude matches that of audio A, preventing audio B from being either masked by or excessively prominent compared to audio A.
Table 4: Rule-based template for generating A+B audio prompts
Condition Probability Sentence Template
$c_{i}<split_{1}$ 30% $P_{B}$, $P_{A}$.
20% $P_{B}$, followed by $P_{A}$.
20% $P_{B}$, then $P_{A}$.
10% with $P_{B}$, $P_{A}$.
10% $P_{B}$, and $P_{A}$.
5% After $P_{B}$, $P_{A}$.
5% $P_{B}$ before $P_{A}$.
$c_{i}>split_{2}$ 30% $P_{A}$, $P_{B}$.
20% $P_{A}$, followed by $P_{B}$.
20% $P_{A}$, then $P_{B}$.
10% $P_{A}$, with $P_{B}$.
10% $P_{A}$, and $P_{B}$.
5% After $P_{A}$, $P_{B}$.
5% $P_{A}$ before $P_{B}$.
$split_{1}\leq c_{i}\leq split_{2}$ 30% $P_{A}$, with $P_{B}$.
30% $P_{A}$, while $P_{B}$.
20% $P_{A}$, $P_{B}$.
20% $P_{A}$, and $P_{B}$.
[…]
Following AUDIT, we reorganize the tuples (A, B, C, A + B, A + C) into the format of (instruction, input audio, output audio), which servers as training data for audio editing tasks. Specifically, (Add, A, A + B) and (Add, A, A + C) constitute two training examples for the add task; (Delete, A + B, A) and (Delete, A + C, A) constitute two training examples for the delete task; (Replace, A + B, A + C) and (Replace, A + C, A + B) constitute two training examples for the replace task.
Appendix D C. Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
Table 5: Number of Samples After Each Round of Selection
Selection Stage Add Delete Replace
Original number 198000 198000 198000
After Round 1 59400 59400 59400
After Round 2 40509 48091 45457
After Round 3 30000 30000 30000
Here, we present the filtering strategy of constructing High-Quality Pseudo-Parallel Audio Editing Dataset. In AppendixC, the construction of the Pseudo-Parallel Audio Editing Dataset is presented, comprising 60,000 training instances for each of the three editing tasks: Addition, Deletion, and Replacement. We employ our proposed model, AuditEval, to evaluate the dataset across three dimensions(Quality, Relevance, and Faithfulness). Subsequently, based on the scores of these three dimensions, we perform three rounds of filtering on the dataset. The filtering criteria are detailed as follows:
  • Stage 1: Faithfulness-based filtering. All samples are sorted by their predicted faithfulness scores in descending order. We retain the top 30% of samples as determined by this ranking.
  • Stage 2: Quality thresholding. Among the retained samples, we further filter out those with predicted quality scores below 3.4.
  • Stage 3: Relevance-based selection. The remaining samples are ranked by predicted relevance scores, and the top 30,000 samples are selected to form the final dataset.
In practice, we use a quality threshold of 3.40, a faithfulness threshold of 3.2156/3.2813/3.2662 and a relevance threshold of 3.1592/3.4844/3.3441. Table5 illustrates the selection process, where the chosen thresholds approximately eliminate half of the data at each round.
Appendix E D. User Interface for Subjective Evaluation
To ensure consistency and ease of use during the manual scoring process, we designed a custom web-based interface for subjective evaluation. As shown in Figure 5, the interface presents the original audio, the edited audio, the original prompt, the edited prompt, and, if available, the edit instruction. Annotators are instructed to score each sample along three dimensions—Quality, Relevance, and Faithfulness—following the same evaluation criteria defined in the AuditScore dataset. This standardized interface streamlines the annotation workflow and helps reduce variability in human ratings.
Figure 5: User Interface for Subjective Evaluation.
Appendix F E. Computational Experiments Details
All experiments were conducted on a single consumer-grade GPU: NVIDIA GeForce RTX 4090. To ensure reproducibility, we fixed all sources of randomness with a global random seed set to 1984.
The AuditEval model was trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$. A learning rate scheduler ReduceLROnPlateau was employed with the following parameters: mode=’min’, factor=0.8, patience=5, and verbose=True.
The training was performed with a batch size of 256 for training and 64 for validation.
The model consists of three independent regression heads corresponding to the Quality, Relevance, and Faithfulness dimensions, trained for different numbers of epochs: Quality: trained for 50 epochs; Relevance: trained for 30 epochs; Faithfulness: trained for 35 epochs.
Appendix G F. Licenses for assets
This work uses publicly available resources. AudioCaps and the audio editing inversion code are released under the MIT License. CLAP pretrained weights are under the CC0-1.0 License. All assets developed in this work—including datasets, checkpoints, training protocols, and code—will be released under the CC BY 4.0 License.
Appendix H G. Ethics Statement
We employed professional annotators via a third-party commercial provider to ensure consistent and reliable subjective evaluations. Annotators were fairly compensated above local minimum wage in accordance with regional labor laws. They received clear instructions on task objectives, scoring criteria (Quality, Relevance, Faithfulness), and example cases prior to annotation. No personal or sensitive information was collected, and no individual-level data beyond task performance was recorded. The provider ensured fair labor practices, contractual compliance, and data confidentiality. As annotators were not research subjects but task workers evaluating system outputs, IRB approval was not required.
Appendix I H. Evaluation scores Across Editing Operations (Add / Delete / Replace)
Algorithm TTA backbone Timesteps System ID Operation Quality score Relevance score Faithfulness score
SDEdit Tango2 75 sys1 Add 3.616 3.206 3.646
Del 3.560 2.466 3.726
Rep 3.686 3.016 3.726
125 sys2 Add 3.656 3.550 3.480
Del 3.602 3.314 3.420
Rep 3.608 3.396 3.226
175 sys3 Add 3.562 3.498 3.108
Del 3.478 3.504 2.896
Rep 3.604 3.508 2.674
Auffusion 75 sys4 Add 3.704 3.356 3.602
Del 3.552 3.122 3.418
Rep 3.550 3.326 3.258
125 sys5 Add 3.518 3.408 3.228
Del 3.412 3.458 2.728
Rep 3.530 3.518 2.700
175 sys6 Add 3.502 3.442 2.826
Del 3.494 3.650 2.724
Rep 3.480 3.500 2.676
DDIM-Inv Tango2 200 sys7 Add 3.522 2.738 3.604
Del 3.358 2.840 3.462
Rep 3.380 2.794 3.402
Auffusion 200 sys8 Add 3.890 3.396 3.454
Del 3.871 3.552 3.420
Rep 3.840 3.392 3.450
ZETA
Tango2 75 sys9 Add 3.798 3.530 3.842
Del 3.620 3.154 3.604
Rep 3.556 3.306 3.478
125 sys10 Add 3.710 3.498 3.588
Del 3.696 3.108 3.506
Rep 3.646 3.456 3.342
175 sys11 Add 3.638 3.518 3.480
Del 3.758 3.658 3.424
Rep 3.568 3.602 3.320
Auffusion 75 sys12 Add 3.622 2.836 3.566
Del 3.666 3.088 3.454
Rep 3.628 3.164 3.428
125 sys13 Add 3.682 3.180 3.482
Del 3.688 3.522 3.460
Rep 3.690 3.592 3.308
175 sys14 Add 3.664 3.302 3.202
Del 3.706 3.546 3.222
Rep 3.636 3.448 3.078
CycleDiffusion Tango2 75 sys15 Add 3.680 3.117 3.528
Del 3.518 2.465 3.690
Rep 3.580 2.960 3.567
125 sys16 Add 3.552 3.442 3.318
Del 3.417 2.940 3.375
Rep 3.365 3.170 3.247
175 sys17 Add 3.240 3.268 3.020
Del 3.283 3.038 3.075
Rep 3.355 3.132 3.067
Auffusion 75 sys18 Add 3.575 3.020 3.583
Del 3.460 2.845 3.517
Rep 3.355 2.900 3.167
125 sys19 Add 3.305 3.035 3.065
Del 3.385 3.145 3.125
Rep 3.390 3.130 2.953
175 sys20 Add 3.177 2.987 2.825
Del 3.318 3.283 2.945
Rep 3.290 3.207 2.907
AP-adapter Audioldm2 - sys21 Add 3.387 2.758 2.675
Del 3.292 2.778 2.835
Rep 3.380 3.035 2.450
Musicgen - - sys22 Add 3.262 2.880 2.673
Del 3.188 3.115 2.793
Rep 3.175 2.910 2.615
Human Edit - - sys23 Add 3.784 3.740 3.940
Del 3.947 4.022 4.037
Rep 3.937 3.891 3.991

### v2

@@
 Each sample is annotated by professional raters on three key aspects of audio editing quality: overall Quality, Relevance to editing intent, and Faithfulness to original features.
-2) Based on this dataset, we train AuditEval, the first model designed for automatic MOS-style scoring tailored to audio editing tasks.
-AuditEval addresses the critical lack of objective evaluation metrics and the prohibitive cost of subjective assessment in this field.
-3) We further leverage AuditEval to evaluate and filter a large amount of synthetically mixed editing pairs, constructing a high-quality pseudo-parallel dataset by selecting the most plausible samples.
-Objective experiments validate the effectiveness of our expert-informed filtering strategy in yielding higher-quality data, while also revealing the limitations of relying solely on objective metrics.
+2) Based on this dataset, we systematically propose AuditEval, a family of automatic MOS-style evaluators tailored for audio editing, covering both SSL-based and LLM-based approaches.
+It addresses the lack of effective objective metrics and the prohibitive cost of subjective evaluation in this field.
+3) We further leverage AuditEval to evaluate and filter a large amount of synthetically mixed editing pairs, mining a high-quality pseudo-parallel subset by selecting the most plausible samples.
+Comprehensive experiments validate that our expert-informed filtering strategy effectively yields higher-quality data, while also exposing the limitations of traditional objective metrics and the advantages of AuditEval.
 The dataset, codes and tools can be found at: https://github.com/NKU-HLT/AuditEval.
-Introduction
-Benefiting from the construction of Text-to-Audio (TTA) and Audio Caption (AC) datasets(Kim et al.
-2019; Gemmeke et al.
-2017; Drossos, Lipping, and Virtanen 2020; Mei et al.
-2024; Piczak 2015; tza 2002), audio generation has made significant strides in recent years.
-Current TTA models—spanning diffusion-based frameworks(Huang et al.
-2023; Liu et al.
-2023, 2024b; Ghosal et al.
-2023; Chen et al.
-2024; Majumder et al.
-2024; Hung et al.
-2024; Xue et al.
-2024) and autoregressive Transformer-based systems(Agostinelli et al.
-2023; Kreuk et al.
+1 Introduction
+Benefiting from the construction of Text-to-Audio (TTA) and Audio Caption (AC) datasets Kim et al.
+(2019); Gemmeke et al.
+(2017); Drossos et al.
+(2020); Mei et al.
+(2024), audio generation has made significant strides in recent years.
+Current TTA models, including diffusion-based frameworks Huang et al.
+(2023); Liu et al.
+(2023, 2024a); Ghosal et al.
+(2023); Chen et al.
+(2024); Majumder et al.
+(2024); Hung et al.
+(2024); Xue et al.
+(2024) and autoregressive transformer-based systems Agostinelli et al.
+(2023); Kreuk et al.
 ; Copet et al.
-2023)—are capable of producing high-quality, diverse, and text-aligned audio outputs based on user prompts.
-However, beyond generation, audio editing remains equally essential and presents more formidable challenges.
-Unlike generation, which involves creating audio from scratch, audio editing requires the system to modify existing audio content according to user-provided textual instructions or target descriptions.
-Unlike generation, which creates audio from scratch, editing modifies existing content based on textual instructions or target descriptions.
+(2023), are capable of producing high-quality, diverse, and text-aligned audio outputs based on textual user prompts.
+However, beyond generation, audio editing remains equally essential but has received comparatively less attention.
+Unlike generation, which creates audio from scratch, audio editing modifies existing audio content based on textual instructions or target descriptions.
 It includes basic event manipulations (e.g., adding, removing, replacing) and more advanced tasks like super-resolution and style transfer.
-Audio editing poses three major challenges: maintaining audio quality, ensuring precise control aligned with user intent, and preserving the unedited portions of the original audio, rather than generating an entirely new sample.
-However, the development of audio editing has been significantly hindered by the lack of high-quality paired datasets—an issue also observed in the vision domain.
-Constructing supervised audio editing datasets is particularly difficult: unlike text-to-audio data, naturally occurring audio pairs that reflect specific editing operations are virtually nonexistent.
-Manual collection is often prohibitively expensive or infeasible—especially for scenarios like replacing gunshots in a crowd with thunder while preserving alignment.
-As a result, most existing methods adopt training-free approaches based on generative diffusion models(Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) and inversion techniques(Song, Meng, and Ermon 2020; Mokady et al.
-2023; Wu and De la Torre 2023; Huberman-Spiegelglas, Kulikov, and Michaeli 2024).
-These methods repurpose the generation and cross-modal alignment capabilities of pretrained TTA models, steering generation toward desired edits by adjusting prompts or conditioning signals(Paissan et al.
-2023; Xu et al.
-2024; Zhang et al.
-2024b; Manor and Michaeli 2024; Jia et al.
-2025; Liu et al.
-2024a; Tsai et al.
-2024).
+Audio editing faces three main challenges: maintaining the overall audio quality, executing edits that accurately reflect user intent, and preserving the unedited portions rather than generating a completely new sample.
+Currently, the development of audio editing has been significantly hindered by the lack of high-quality paired datasets.
+This has also been observed in the image editing domain.
+Constructing supervised audio editing datasets is particularly difficult.
+Unlike TTA data, naturally occurring audio pairs that reflect specific editing operations are virtually nonexistent.
+Manual collection is often prohibitively expensive or infeasible, especially for scenarios like replacing gunshots in a crowd with thunder while preserving alignment.
+As a result, most existing methods adopt training-free approaches based on generative diffusion models Ho et al.
+(2020); Song et al.
+(2020) and inversion techniques Song et al.
+(2020); Mokady et al.
+(2023); Wu and De la Torre (2023); Huberman-Spiegelglas et al.
+(2024).
+These methods repurpose the generation and cross-modal alignment capabilities of pretrained TTA models, steering generation toward desired edits by adjusting prompts or conditioning signals Paissan et al.
+(2023); Xu et al.
+(2024); Zhang et al.
+(2024b); Manor and Michaeli (2024); Jia et al.
+(2025); Liu et al.
+(2024b); Tsai et al.
+(2024).
 However, they heavily depend on the underlying TTA model and often lack robustness due to the absence of task-specific supervision.
 As an alternative to training-free approaches, supervised frameworks have also been explored by training models on explicitly constructed editing datasets.
-Audit(Wang et al.
-2023b) represents the first supervised audio editing framework, introducing a triplet-based dataset and training a dedicated latent diffusion model (LDM).
+Audit Wang et al.
+(2023) represents the first supervised audio editing framework, introducing a triplet-based dataset and training a dedicated latent diffusion model (LDM).
 To date, it remains the only task-specific paired dataset for audio editing.
-This dataset is created through audio mixing and combinatorial pairing strategies, leveraging the compositional nature of audio—unlike images, audio signals can be naturally combined to simulate coherent edits.
-A similar construction paradigm is adopted in InstructME(Han et al.
-2023), which applies mixing-based strategies to generate supervision signals for music editing tasks.
-Meanwhile, MusicGen(Copet et al.
-2023) and Instruct-MusicGen(Zhang et al.
-2024a) explore autoregressive approaches based on discrete audio-text tokens, supporting tasks such as continuation, transformation, and editing.
-WavCraft(lia 2024) further investigates instruction-based editing by employing large language models (LLMs) to interpret textual commands and orchestrate external audio and speech models for modification.
-Despite recent progress, there is still no unified standard for evaluating audio editing tasks.
-The absence of a widely accepted, open-source benchmark specifically designed for editing has led to fragmented evaluation protocols and a lack of consistent objective metrics.
-Currently, the most commonly used metric is the CLAP Score, which measures similarity between the edited audio and the target prompt in the embedding space, leveraging the CLAP model(Elizalde et al.
-2023) for cross-modal comparison without requiring ground-truth audio.
-To assess preservation of unedited regions, some works(Paissan et al.
-2023; Zhang et al.
-2024b; Jia et al.
-2025; Liu et al.
-2024a) use the original audio as a reference and compute similarity between the original and edited versions.
-Others(Xu et al.
-2024; Zhang et al.
-2024b) construct pseudo-paired datasets via audio mixing—following a strategy similar to Audit and InstructME—and evaluate outputs by comparing them to mixed signals using conventional TTA metrics such as Fréchet Distance (FD), Fréchet Audio Distance (FAD)(Roblek et al.
-2019), and KL divergence.
-However, such evaluation paradigms—relying on synthetic mixtures as ground truth—often fall short.
-Since high-quality audio edits may vary significantly and are not strictly aligned with a single reference, these metrics can fail to reliably capture true editing quality or model controllability.
+This dataset is created through audio mixing and combinatorial pairing strategies.
+It leverages the compositional nature of audio, which unlike images, can be naturally mixed to simulate coherent edits.
+A similar construction paradigm is adopted in InstructME Han et al.
+(2023), which applies mixing-based strategies to generate supervision signals for music editing tasks.
+Meanwhile, MusicGen Copet et al.
+(2023) and Instruct-MusicGen Zhang et al.
+(2024a) explore autoregressive approaches based on discrete audio-text tokens, supporting tasks such as continuation, transformation, and editing.
+WavCraft Liang et al.
+further investigates instruction-based editing by employing large language models (LLMs) to interpret textual commands and call external audio and speech models for modification.
+Despite these progress, there is still no unified standard for evaluating audio editing tasks.
+The absence of a widely accepted, open-source benchmark specifically designed for editing has led to fragmented evaluation protocols and a lack of reliable objective metrics.
+Currently, the most commonly used metric is the CLAP_Score, which measures similarity between the edited audio and the target prompt in the embedding space, leveraging the CLAP model Elizalde et al.
+(2023) for cross-modal comparison without requiring ground-truth audio.
+To assess preservation of unedited regions, some works Paissan et al.
+(2023); Zhang et al.
+(2024b); Jia et al.
+(2025); Liu et al.
+(2024b) use the original audio as a reference and compute similarity between the original and edited versions.
+Others Xu et al.
+(2024); Zhang et al.
+(2024b) construct pseudo-paired datasets via audio mixing, following a strategy similar to that of Audit, and evaluate the generated outputs by comparing them with the mixed signals using conventional TTA metrics such as Fréchet Distance (FD), Fréchet Audio Distance (FAD) Roblek et al.
+(2019) and KL divergence.
+However, such evaluation paradigms, which rely on synthetic mixtures as ground truth, often fall short.
+Because high-quality audio edits can vary substantially and are not necessarily aligned with a single reference.
 As a result, subjective human evaluation remains the most reliable method for assessing audio editing quality.
-Current subjective evaluation protocols typically consider three aspects: overall audio Quality(Zhang et al.
-2024b; Jia et al.
-2025), Relevance to the editing intent(Paissan et al.
-2023; Xu et al.
-2024; Zhang et al.
-2024b; Manor and Michaeli 2024; Jia et al.
-2025; Liu et al.
-2024a), and Faithfulness to the original (unedited) content(Paissan et al.
-2023; Xu et al.
-2024; Zhang et al.
-2024b; Manor and Michaeli 2024; Jia et al.
-2025; Liu et al.
-2024a).
+Current subjective evaluation protocols typically consider three aspects: overall audio Quality Zhang et al.
+(2024b); Jia et al.
+(2025), Relevance to the editing intent Paissan et al.
+(2023); Xu et al.
+(2024); Zhang et al.
+(2024b); Manor and Michaeli (2024); Jia et al.
+(2025); Liu et al.
+(2024b), and Faithfulness to the original (unedited) content Paissan et al.
+(2023); Xu et al.
+(2024); Zhang et al.
+(2024b); Manor and Michaeli (2024); Jia et al.
+(2025); Liu et al.
+(2024b).
 These criteria provide an intuitive and comprehensive assessment of audio editing performance.
-Nevertheless, compared to pure generation tasks such as Text-to-Audio (TTA), Text-to-Speech (TTS), or Text-to-Music (TTM), audio editing is inherently more complex, requiring deeper understanding and recognition of nuanced content.
+Nevertheless, compared to pure generation tasks such as TTA, Text-to-Speech (TTS), or Text-to-Music (TTM), audio editing is inherently more complex, requiring deeper understanding and recognition of nuanced paired contents.
 This results in increased costs for annotator training, longer evaluation durations, and greater human labor requirements in practice.
-In light of these challenges, developing automatic evaluation methods and building a dedicated evaluation system for audio editing is both meaningful and valuable for advancing research in this field.
-In this paper, we aim to bridge the evaluation gap in audio editing by introducing AuditEval, a scalable automatic MOS-style assessment framework, and subsequently leveraging it to build a high-quality pseudo-parallel dataset.
+To these challenges, developing automatic Mean Opinion Score (MOS) prediction methods and frameworks for audio editing is extremely important and meaningful for advancing research in this field.
+In this paper, we aim to bridge the data and evaluation gaps in audio editing by introducing AuditEval, a scalable automatic MOS-style assessment framework, and subsequently leveraging it to help mining high-quality pseudo-parallel audio editing data.
 Our contributions are threefold:
-  • Subjective Audio Editing Evaluation Dataset Construction.
-We construct AuditScore, the first comprehensive benchmark for subjective audio editing evaluation, consisting of over 6,300 edited samples from 7+ representative frameworks and 23+ system variants.
-Each sample is annotated by professional raters on three key aspects: Quality, Relevance, and Faithfulness.
-  • Automatic MOS Prediction.
-Based on AuditScore, we train AuditEval, the first model tailored for automatic MOS-style scoring in audio editing.
-It effectively mitigates the lack of objective metrics and the high cost of human evaluation.
-  • High-Quality Pseudo-Parallel Audio Editing Dataset Construction.
-We further utilize AuditEval to assess and filter a large volume of synthetically mixed editing pairs, selecting only the most plausible samples to construct a high-quality pseudo-parallel dataset.
-Objective experiments validate the effectiveness of this expert-informed filtering strategy in producing superior-quality data, while also highlighting the shortcomings of relying exclusively on objective metrics.
-Related Work
-Training-free Audio Editing Method
+  • Evaluation Dataset Construction.
+We construct AuditScore, the first comprehensive dataset for subjective audio editing evaluation, consisting of over 6,300 edited pairs from 7+ representative frameworks and 23+ system variants.
+Each sample is annotated by professional raters from three key aspects: Quality, Relevance, and Faithfulness.
+  • Automatic MOS-style Evaluation.
+Based on AuditScore, we design and systematically study two paradigms for automatic MOS-style assessment in audio editing: a Self-Supervised Learning (SSL)-based evaluator (AuditEval-ssl) and a Large Language Model (LLM)-based evaluator (AuditEval-llm).
+We aim to provide scalable and cost-effective alternatives to expensive human evaluation.
+  • High-Quality Data Mining.
+We further utilize AuditEval to assess and filter a large amount of synthetically mixed editing pairs, selecting only the most plausible samples to form a high-quality pseudo-parallel subset.
+Comprehensive experiments validate the effectiveness of this expert-informed filtering strategy in producing superior-quality data, while also highlighting the shortcomings of objective metrics and the advantages of AuditEval in audio editing task.
+2 Related Work
+2.1 Training-free Audio Editing Method
 Training-free audio editing methods exploit pre-trained TTA diffusion models to perform editing without the need for task-specific fine-tuning.
-These approaches typically begin by inverting the input audio into the noise space of the diffusion model—that is, mapping from $x_{0}$ (the observed audio) to $x_{T}$ (a noisy latent representation).
+These approaches typically begin by inverting the input audio into the noise space of the diffusion model, that is, mapping from $x_{0}$ (the observed audio) to $x_{T}$ (a noisy latent representation).
 The resulting latent $x_{T}$ preserves the semantic and structural features of the original audio while providing a flexible representation for guided generation.
@@
 $\boldsymbol{s}_{\theta}(\mathbf{x_{t}},t)+\sqrt{\textstyle\sigma^{2}(t)\!-\!\sigma^{2}(t+\Delta t)}\,\mathbf{z}$
-DDIM-based approaches perform deterministic inversion of the diffusion process by exploiting the approximate reversibility of the underlying ODE with sufficiently small step sizes (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Mokady et al.
-2023).
+DDIM-based approaches perform deterministic inversion of the diffusion process by exploiting the approximate reversibility of the underlying ODE with sufficiently small step sizes Ho et al.
+(2020); Song et al.
+(2020); Mokady et al.
+(2023).
 This allows for editing by first reversing the generation trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}T$ (Eq.
@@
 3), thereby achieving semantically guided edits.
-Building upon DDIM inversion, several studies (Xu et al.
-2024; Zhang et al.
-2024b; Jia et al.
-2025; Liu et al.
-2024a) incorporate enhancements such as attention map manipulation and null-text optimization, demonstrating effective results in controllable audio editing.
+Building upon DDIM inversion, several studies  Xu et al.
+(2024); Zhang et al.
+(2024b); Jia et al.
+(2025) incorporate enhancements such as attention manipulation and null-text optimization, achieving effective results in controllable audio editing.
 $x_{t+1}=\textstyle\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}x_{t}+\left(\sqrt{\frac{1-\alpha_{t+1}}{\alpha_{t+1}}}\!-\!\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}\right)\cdot\varepsilon_{\theta}(x_{t},t,\mathcal{P}_{\boldsymbol{ori}})$ (2)
 $x_{t-1}=\textstyle\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}x_{t}+\left(\sqrt{\frac{1-\alpha_{t-1}}{\alpha_{t-1}}}\!-\!\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}\right)\cdot\varepsilon_{\theta}(x_{t},t,\mathcal{P}_{\boldsymbol{tar}})$ (3)
-DDPM-based approaches include CycleDiffusion(Wu and De la Torre 2023) and Edit-Friendly DDPM(Huberman-Spiegelglas, Kulikov, and Michaeli 2024).
-This kind of method leverages the stochastic nature of the DDPM sampling process—specifically, the stochastic noise injected at each diffusion step—to preserve the characteristics of the original input.
+DDPM-based approaches include CycleDiffusion Wu and De la Torre (2023) and Edit-Friendly DDPM Huberman-Spiegelglas et al.
+(2024).
+This kind of method leverages the stochastic nature of the DDPM sampling process, specifically, the stochastic noise injected at each diffusion step to preserve the characteristics of the original input.
 By adding noise to the original input and storing intermediate latent states at each step, the model can compute the corresponding stochastic noise components (Eq.4, $\mu_{t}\left(\mathrm{x}_{t}\right)$ is the calculated expectation of ${x}_{t}$, $\sigma_{t}$ is the hyperparameter), which are then reused during the denoising process under the modified prompt to retain original content.
-Edit-Friendly DDPM further improves by altering the noise injection mechanism so that the noise added at each timestep is independent, thereby enhancing editing controllability and performance(Huberman-Spiegelglas, Kulikov, and Michaeli 2024).
-ZETA Editing(Manor and Michaeli 2024) is the first known method to apply DDPM inversion to audio editing, demonstrating superior results compared to SDEdit- and DDIM-inversion-based methods.
+Edit-Friendly DDPM further improves by altering the noise injection mechanism so that the noise added at each timestep is independent, thereby enhancing editing controllability and performance.
+ZETA Editing Manor and Michaeli (2024) is the first known method to apply DDPM inversion to audio editing, demonstrating superior results compared to SDEdit- and DDIM-inversion-based methods.
 $\mathbf{z}_{t}=\left(\mathrm{x}_{t-1}-\mu_{t}\left(\mathrm{x}_{t}\right)\right)/\sigma_{t},\quad t=T,\ldots,1$ (4)
-Supervised Audio Editing Models
+2.2 Supervised Audio Editing Models
 In contrast to training-free methods, supervised audio editing approaches rely on explicitly constructed editing datasets and can be broadly divided into non-autoregressive and autoregressive paradigms.
 Models such as Audit and InstructME employ latent diffusion models conditioned on audio-text multimodal inputs.
-They utilize specialized conditioning mechanisms and position-aware loss functions to enable end-to-end training on pseudo-labeled editing pairs or instrument transfer datasets, achieving effective editing performance (Wang et al.
-2023b; Han et al.
-2023).
-On the other hand, autoregressive approaches generate edited audio by conditioning on discrete audio-text representations of the original content and edit instructions as context.
-Representative models like MusicGen and Instruction-MusicGen adopt transformer-based decoders to sequentially synthesize edited audio (Copet et al.
-2023; Zhang et al.
-2024a).
+They utilize specialized conditioning mechanisms and position-aware loss functions to enable end-to-end text-controlled audio editing Wang et al.
+(2023); Han et al.
+(2023).
+On the other hand, autoregressive approaches generate edited audio by conditioning on discrete representations of the original audio and textual edit instruction as context.
+Representative model like MusicGen adopt transformer-based decoders to sequentially synthesize edited audio  Copet et al.
+(2023); Zhang et al.
+(2024a).
 While these models benefit from strong language-audio alignment and high generation quality, they often struggle with temporal inconsistencies between the original and edited audio.
-Automatic Evaluation for Speech, Music and Audio
+2.3 Auto-Evaluation for Speech, Music and Audio
 Automatic subjective evaluation plays a vital role in assessing AI-generated outputs by simulating human perception in a scalable and repeatable manner.
-Compared to purely objective metrics, it provides a more intuitive and comprehensive measure of perceptual quality.
-A well-designed evaluator not only aligns with human preferences but also guides the development of future generative models.
-In the speech domain, automatic Mean Opinion Score (MOS) prediction has been widely adopted to evaluate tasks such as Text-to-Speech (Cooper et al.
-2022; Wang et al.
-2023a, 2024), Text-to-Voice (Tang et al.
-2024), and speech enhancement(Reddy, Gopal, and Cutler 2021; Mittag et al.
-2021; Chen et al.
-2022).
-These models have shown promising results and effectively complement traditional objective metrics.
-In contrast, automatic evaluation for audio and music remains more challenging due to the greater content diversity—from natural sound events to complex musical structures—and the inherently looser alignment between audio and textual descriptions compared to the tighter correspondence in speech-transcription pairs.
-To address these challenges, MusicEval(Liu et al.
-2025) is the first known framework to provide automatic scoring for music generation.
-It constructs a high-quality evaluation dataset and trains scoring models to assess both overall quality and text relevance, achieving strong correlation with human judgments.
-Meanwhile, Meta’s Audiobox Aesthetics(Tjandra et al.
-2025) proposes a unified framework for speech, music, and audio evaluation, defining four core aesthetic dimensions that collectively provide a comprehensive and structured standard for automatic audio assessment from an aesthetic perspective.
-Data Collection
-Basic Information of AuditScore
-Table 1: Subjective evaluation result. Detailed distribution of difference editing types is presented in AppendixH I.
-Editing
-Model or Method
-Generative
-TTA Backbone
-Timesteps
-System
-ID
-Mean Scores
-Quality
-Relevance
-Faithfulness
-SDEdit Tango2 75 sys1 3.621 2.896 3.699
-125 sys2 3.622 3.420 3.375
-175 sys3 3.548 3.503 2.893
-Auffusion 75 sys4 3.602 3.268 3.426
-125 sys5 3.487 3.461 2.885
-175 sys6 3.492 3.531 2.742
-DDIM Inversion Tango2 200 sys7 3.420 2.791 3.489
-Auffusion 200 sys8 3.867 3.447 3.441
-Edit-friendly
-DDPM Inversion
-Tango2 75 sys9 3.658 3.330 3.641
-125 sys10 3.684 3.354 3.479
-175 sys11 3.655 3.593 3.408
-Auffusion 75 sys12 3.639 3.029 3.483
-125 sys13 3.687 3.431 3.417
-175 sys14 3.669 3.432 3.167
-CycleDiffusion
-DDPM Inversion
-Tango2 75 sys15 3.592 2.848 3.595
-125 sys16 3.445 3.184 3.313
-175 sys17 3.293 3.146 3.054
-Auffusion 75 sys18 3.463 2.922 3.423
-125 sys19 3.360 3.103 3.047
-175 sys20 3.262 3.159 2.893
-AP-adapter Audioldm2 - sys21 3.353 2.857 2.653
-Musicgen - - sys22 3.208 2.968 2.693
-Human Edit - - sys23 3.883 3.878 3.988
+Compared to objective metrics, it provides a more intuitive measure of perceptual quality.
+A well-designed evaluator not only aligns with human preferences but also guides the development of future models.
+In the speech domain, automatic Mean Opinion Score (MOS) prediction has been widely adopted to evaluate tasks such as Text-to-Speech Cooper et al.
+(2022); Wang and et al.
+(2023), Text-to-Voice Tang et al.
+(2024), and speech enhancement Reddy et al.
+(2021); Mittag et al.
+(2021); Chen et al.
+(2022).
+These models have shown promising results and effectively complement objective metrics.
+In contrast, automatic evaluation for audio and music is more challenging.
+The content is highly diverse, ranging from natural sound events to complex musical structures.
+Moreover, the alignment between audio and textual descriptions is inherently looser than the tight correspondence found in speech transcription.
+To address these challenges, MusicEval Liu et al.
+(2025) and AudioEval Wang et al.
+(2025) construct high-quality evaluation datasets and train scoring models to assess both overall quality and text relevance, achieving strong correlations with human judgments.
+Meanwhile, Meta’s Audiobox Tjandra et al.
+(2025) proposes a unified framework for speech, music, and audio evaluation, defining four core aesthetic dimensions that collectively provide a comprehensive and structured standard for automatic audio assessment from an aesthetic perspective.
+3 Dataset Construction
+3.1 Basic Information of AuditScore
+Editing Model or Method Generative TTA Backbone Timesteps System IDs Average Editing Scores
+Quality Relevance Faithfulness
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Subjective evaluation results across various systems. Detailed distribution of different editing types is presented in Appendix I.
 AuditScore is a human-annotated dataset comprising 6,360 pairs of original and edited audio samples, totaling 35.3 hours in duration.
-These samples are generated using 23 distinct system configurations across seven representative audio editing approaches: SDEdit (Meng et al.
-), DDIM (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020), Edit-Friendly DDPM (Huberman-Spiegelglas, Kulikov, and Michaeli 2024; Manor and Michaeli 2024), CycleDiffusion (Wu and De la Torre 2023), AP-adapter (Tsai et al.
-2024), MusicGen (Copet et al.
-2023), and Human Edit(manual editing).
-For diffusion-based methods (SDEdit, DDIM, Edit-Friendly DDPM, and CycleDiffusion), we adopt two state-of-the-art generative backbones—Tango2 (Majumder et al.
-2024) and Auffusion (Xue et al.
-2024)—to synthesize the edited audio.
+These samples are generated using 23 distinct system configurations across seven representative audio editing approaches: SDEdit, DDIM, Edit-Friendly DDPM, CycleDiffusion, AP-adapter, MusicGen, and Human Edit(manual editing).
+For diffusion-based methods (SDEdit, DDIM, Edit-Friendly DDPM, and CycleDiffusion), we adopt two state-of-the-art generative backbones Tango2  Majumder et al.
+(2024) and Auffusion  Xue et al.
+(2024) to synthesize the edited audio.
 To further promote diversity and enhance model generalization, each configuration is instantiated with three inversion timesteps (75, 125, and 175 out of 200), allowing finer control over the degree of editing.
 This setup is specifically applied to SDEdit, Edit-Friendly DDPM, and CycleDiffusion to explore a wider range of editing intensities and semantic variations.
-To construct a representative and diverse dataset, we began by selecting original audio samples and corresponding captions from the AudioCaps dataset (Kim et al.
-2019).
-Captions were clustered into 100 semantic groups using GTE-Large text embeddings (Li et al.
-2023), from which 2–3 audio samples were manually selected per cluster.
-For each selected audio, we crafted three types of editing instructions—Addition, Deletion, and Replacement—yielding a shared pool of 240 instructions across all systems.
-Additionally, several representative systems were assigned 30 unique instructions to further increase variation in editing scenarios.
-In total, AuditScore covers over 200 distinct audio event types, with editing operations introducing more than 150 novel events.
-The rich variety of methods, model architectures, editing parameter settings, and content diversity enables AuditScore to capture a wide spectrum of editing behaviors, audio qualities, and modeling challenges—making it a comprehensive and robust benchmark for training and evaluating audio editing evaluation models.
+To construct a representative and diverse dataset, we first select original audio samples and corresponding captions from AudioCaps.
+The captions are clustered into 100 semantic groups using GTE-large Li et al.
+(2023) text embeddings, from which 2 to 3 audio samples are manually selected per cluster.
+For each selected audio, we design three types of editing instructions, including Addition, Deletion, and Replacement, resulting in a shared pool of 240 instructions across all systems.
+In addition, each representative systems are assigned 30 unique instructions to further increase the diversity of editing scenarios.
+Overall, AuditScore covers more than 200 distinct audio event types, with editing operations introducing over 150 novel events.
+The diversity in algorithms, model architectures, editing parameters, and audio content enables AuditScore to capture a wide range of editing behaviors and quality variations, making it a robust source of expert-informed knowledge for training and evaluating automatic audio editing evaluation models.
 To assess the edited audio samples generated by these 23 systems, we conducted a large-scale subjective listening study involving five domain experts with extensive experience in audio dataset curation and perceptual evaluation.
 Each expert independently rated the editing audio pairs along three key perceptual dimensions: Quality, Relevance, and Faithfulness.
-The final rating for each system was computed as the average of expert scores across all three dimensions.
+The final rating for each utterance was computed as the average of five expert scores across individual dimensions.
 A comprehensive summary of system IDs and their evaluation results is presented in Table 1.
-(a) Quality (Distribution)
-(b) Relevance (Distribution)
-(c) Faithfulness (Distribution)
-(d) Quality (Heatmap)
-(e) Relevance (Heatmap)
-(f) Faithfulness (Heatmap)
-Figure 1: Comparison of quality, relevance, and faithfulness across all systems and annotation types. Top: score distribution across samples. Bottom: system-level heatmaps from AuditScore.
-Scoring Protocol and Analysis
+Figure 1: Distributions of AuditScore annotations (left) and AuditEval predictions (right) at both the system and utterance levels.
+3.2 Scoring Protocol and Analysis
 To comprehensively evaluate audio editing performance, we adopt a three-dimensional scoring protocol encompassing Quality, Relevance, and Faithfulness.
-These dimensions are informed by established perceptual evaluation methodologies—such as MOS (Mean Opinion Score), CMOS (Comparison MOS)(Loizou 2011), and semantic alignment assessments—while tailored specifically for the nuances of audio editing.
+These dimensions are informed by established perceptual evaluation methodologies, such as MOS (Mean Opinion Score), CMOS (Comparison MOS) Loizou (2011), and semantic alignment assessments, and are further tailored to capture the specific nuances of audio editing.
 Detailed annotation guidelines were developed to ensure consistency across expert raters while preserving flexibility for nuanced perceptual judgments.
-The scoring scales are defined as follows(Additional scoring rubrics and details can be found in AppendixA B):
+The scoring scales are as follows(details can be found at Appendix A):
   • Quality —Assesses the overall semantic correctness and acoustic integrity of the edited audio: (1) About the Same / Even Better, (2) Slightly Worse, (3) Noticeably Worse, (4) Much Worse, (5) Severely Degraded.
@@
   • Faithfulness —Evaluates the preservation of unedited content from the original audio in terms of rhythm, timing, loudness, and overall style: (1) Perfectly Faithful, (2) Mostly Faithful, (3) Partially Faithful, (4) Minimally Faithful, (5) Not Faithful.
-Figure 1a–c show the distribution of scores across the three dimensions.
+Figure 1 the leftmost column, shows the distribution of scores across the three dimensions.
 The scores for Quality, Relevance, and Faithfulness each approximately follow a normal distribution, indicating that our scoring criteria provide a balanced and well-calibrated assessment scale.
-Notably, quality scores are slightly skewed toward higher values, reflecting that most editing systems tend to preserve audio quality easily, while still facing challenges in precisely executing editing instructions and maintaining content fidelity.
-Figure 1b–d further visualize performance across different systems.
+Notably, quality scores are slightly skewed toward higher values, reflecting that most editing systems tend to maintain audio quality easily, while still facing challenges in precisely executing editing instructions and preserving unedited content.
+Figure 1 the middle-left column, further visualizes performance across different systems.
 Notably, Human Editing consistently achieves the highest ratings in all dimensions, establishing an upper bound for automated models.
-Interestingly, some systems (e.g., sys3, sys5, and sys6) achieve high relevance scores, indicating strong alignment with the editing instructions, yet score lower in faithfulness—highlighting a common trade-off in current models between effective editing and content preservation.
+Interestingly, some systems (e.g., sys3, sys5, and sys6) achieve high relevance scores, indicating strong alignment with the editing instructions, yet score lower in faithfulness, highlighting a common trade-off in current models between effective editing and content preservation.
 This analysis underscores the remaining gap between human and machine editing and suggests a clear direction for future model improvement.
-Figure 2: The overall architecture of AuditEval.
-Table 2: Performance of the proposed model on the AuditScore dataset at system and utterance levels across three dimensions.
-Dimension System-level Utterance-level
-MSE↓ LCC↑ SRCC↑ KATU↑ MSE↓ LCC↑ SRCC↑ KATU↑
-Quality Score 0.0595 0.6101 0.7004 0.5516 0.1474 0.3318 0.3306 0.2462
-Relevance Score 0.0962 0.6196 0.6472 0.4783 0.2963 0.4105 0.3991 0.2886
-Faithfulness Score 0.0726 0.8460 0.8809 0.7024 0.2217 0.5327 0.5171 0.3791
-Table 3: Comparison of pseudo-parallel data quality before and after filtering.
-Objective metrics include CLAPt (cross-modal similarity between edited audio and target text), CLAPa(similarity between edited audio and target audio), IS, FAD, FD, and KL divergence (all measuring distributional similarity between edited and original audio).
-Subjective metrics are predicted human preference scores from AuditEval, used to compare perceived quality and reveal discrepancies with objective measures.
-Dataset Objective Subjective (Automatic)
-CLAPt(%)↑ CLAPa(%)↑ IS↑ FAD↓ FD↓ KL↓ Quality Relevance Faithfullness
-Random1 48.70 82.73 12.23 0.2859 3.207 0.6913 3.365 3.281 3.114
-Random2 48.47 82.69 12.23 0.2816 3.197 0.6892 3.365 3.283 3.114
-Random3 48.76 82.73 12.20 0.2830 3.205 0.6867 3.366 3.281 3.114
-Filtered 49.43 82.76 12.61 0.4587 4.356 0.6783 3.626 3.694 3.493
-Method
-AuditEval: An Automatic Evaluation Model
-Based on our newly curated AuditScore dataset, we develop AuditEval—an automatic MOS-style evaluation model for audio editing built upon the MusicEval framework.
-AuditEval processes two groups of inputs: the original audio paired with its edited version, and the corresponding textual descriptions of the audio before and after editing, respectively.
-It outputs three scalar scores reflecting the key evaluation dimensions: Quality, Relevance, and Faithfulness.
-To account for the different learning complexities across these dimensions, AuditEval employs three independently trained decoding heads, each specialized for predicting one specific score.
-This simple yet effective design represents a minimal adaptation of the MusicEval framework (see Figure 2), with modest architectural changes and some optimization efforts, while still achieving balanced and reliable predictions across all evaluation aspects.
-Pseudo-parallel Data Filtering
-Due to the lack of well-established supervised datasets for audio editing, we construct a large-scale (500,000 pairs, 2700h) Pseudo-parallel Audio Editing Dataset following the methodologies of AUDIT (Wang et al.
-2023b) and InstructME (Han et al.
-2023), and further enhance it using our automatic scoring model, AuditEval.
-The dataset consists of synthetic audio tuples (A, B, C, A+B, A+C), where Audio A is sampled from clustered AudioCaps (Kim et al.
-2019) descriptions, and B/C are single-event clips paired with semantically similar prompts.
+4 Method
+Figure 2: Model designs of AuditEval-llm and AuditEval-ssl, along with the corresponding training modules and inference workflows.
+Dimensions Models System-level Prediction Utterance-level Prediction
+[…]
+Quality Obj-Metric 8.7086 0.2277 0.2470 0.1858 - - - -
+AuditEval-ssl 0.0595 0.6101 0.7004 0.5516 0.1474 0.3318 0.3306 0.2462
+AuditEval-llm 0.0198 0.5814 0.5523 0.3897 0.0999 0.3840 0.3967 0.2922
+Relevance Obj-Metric 8.7884 0.2141 0.1779 0.1225 9.0698 0.1344 0.0812 0.0574
+AuditEval-ssl 0.0962 0.6196 0.6472 0.4783 0.2963 0.4105 0.3991 0.2886
+AuditEval-llm 0.0328 0.7620 0.8142 0.6443 0.2189 0.4766 0.4049 0.2929
+Faithfulness Obj-Metric 6.7973 0.8790 0.8843 0.7024 6.9466 0.4981 0.5038 0.3611
+AuditEval-ssl 0.0726 0.8460 0.8809 0.7024 0.2217 0.5327 0.5171 0.3791
+AuditEval-llm 0.0285 0.9133 0.8952 0.7579 0.1313 0.7213 0.6898 0.5300
+Table 2: Performance comparison of score prediction at both system and utterance levels.
+denotes the Inception Score (IS), distribution-level metric for Quality, denotes the cross-modal CLAP score for Relevance, and denotes the audio–audio CLAP score for Faithfulness.
+We present AuditEval, a scalable automatic MOS-style evaluation framework for audio editing that explores two complementary paradigms: an SSL-based predictor (AuditEval-ssl) and an LLM-based evaluator (AuditEval-llm), both producing aspect-wise scores for Quality, Relevance, and Faithfulness.
+4.1 AuditEval-SSL
+AuditEval-ssl is designed to automatically predict MOS-style scores for audio editing results by leveraging self-supervised audio representations.
+Given an original audio, an edited audio, and their corresponding original and target descriptions, the model first extracts high-level acoustic features and temporal cues from the audio using a pretrained SSL backbone (we adopt ms-clap, which can automatically align audio and text semantics), and maps these representations to subjective scores along three evaluation dimensions.
+The core idea of AuditEval-ssl is to transfer the general-purpose audio understanding capability of SSL models to the audio editing evaluation setting, enabling robust and scalable scoring without relying on paired ground-truth edited audio.
+4.2 AuditEval-LLM
+AuditEval-llm explores an alternative evaluation paradigm in which a large language model serves as an evaluator by reasoning over multimodal evidence.
+Instead of directly regressing scores from latent embeddings, we convert the input audios together with their textual captions into a structured textual evidence set (details can be found at Appendix H) and prompt the LLM to produce MOS-style ratings.
+In terms of model design, we adopt the advanced Large Audio Language Model Qwen2.5-Omni-7B.
+Since the target scores are continuous float values obtained by averaging five independent human ratings, we move beyond conventional dialog-style discrete scoring in a {1–5} set.
+Specifically, we retain the Qwen2.5-Omni Thinker module and design three post regression heads to predict continuous scores for the three evaluation dimensions.
+This design leverages the instruction-following, cross-modal understanding, and judgment reasoning capabilities of LLMs for audio editing evaluation.
+4.3 Pseudo-parallel Data Filtering
+Due to the lack of well-established supervised datasets for audio editing, we construct a large-scalePseudo-parallel Audio Editing Dataset following the methodologies of AUDIT Wang et al.
+(2023) and InstructME Han et al.
+(2023), and further enhance it using our automatic scoring model, AuditEval.
+The dataset consists of synthetic audio tuples (A, B, C, A+B, A+C), where Audio A is sampled from clustered AudioCaps Kim et al.
+(2019) descriptions, and B/C are single-event clips paired with semantically similar prompts.
 These are reorganized into instruction-based triplets for Addition, Deletion, and Replacement tasks.
 A rule-based strategy is used to mix the audio and construct corresponding textual instructions, aiming to maintain acoustic coherence and semantic plausibility.
-Details are provided in AppendixB C.
+Details are provided in Appendix B.
 However, this kind of rule-based concatenation introduces several limitations: (1) the quality and prompt relevance of individual audio elements are inconsistent; (2) the insertion positions in the background audio are often suboptimal—even when attempting to insert at quieter regions, the resulting mixtures can sound unnatural and make it difficult to distinguish the editing change; (3) the randomly selected A/B/C combinations can be semantically implausible—for example, inserting the sound of paper rustling into a battlefield recording, or replacing male speech with a toilet flushing sound, etc.
-To address these issues, we apply AuditEval to automatically evaluate and filter synthesized audio editing pairs based on scoring criteria informed by AuditScore.
-Because AuditScore covers almost all mainstream editing operations including mixing, this filtering is expected to improve data quality and semantic plausibility, advancing toward a practical and scalable supervised dataset (see AppendixC D).
-Experiments
-Dataset and Metrics
-For the automatic evaluation of audio editing, we employ our proposed AuditEval model and train it on the curated AuditScore dataset using a fixed 9:1 train-test split.
-The dataset will be publicly released to facilitate reproducibility.
-We report the Mean Squared Error (MSE) at both the utterance and system levels to quantify prediction accuracy.
-To evaluate consistency with human judgments, we additionally report the Linear Correlation Coefficient (LCC), Spearman’s Rank Correlation Coefficient (SRCC), and Kendall’s Tau Correlation Coefficient (KTAU)(Cooper et al.
-2022), which capture linear and rank-based relationships between predicted scores and ground-truth annotations.
-For the task of selecting high-quality pseudo-parallel audio editing data, we conduct exploratory experiments with objective evaluations to compare the audio quality before and after filtering.
-Referring to typical training data scales, we retain the top 15% of data for comparison.
-Specifically, we adopt FAD, FD, and KL, following the methodology of (Liu et al.
-2023), to assess the distributional similarity between audio sets.
-Additionally, we use the CLAP Score to evaluate cross-modal alignment between audio and corresponding descriptions.
-(a) Quality
-(b) Relevance
-(c) Faithfulness
-Figure 3: System-level comparison between AuditEval predictions and human ratings across three evaluation aspects.
-Results and Analysis
-We first evaluate the effectiveness of our automatic evaluation model, AuditEval, which is trained on the curated AuditScore dataset.
-Table 2 presents the evaluation results at both the system and utterance levels.
-At the system level, the mean squared error (MSE) between AuditEval predictions and human ratings remains impressively low—below 0.1 on a 1–5 scale—across all evaluation dimensions.
-Meanwhile, the linear correlation coefficient (LCC) exceeds 0.6 in every case.
-Notably, Faithfulness achieves an LCC of 0.84, indicating strong agreement with human judgments.
-As illustrated in Figure 3, AuditEval effectively ranks different models according to their performance, particularly in terms of executing editing intentions (Relevance) and preserving unedited content (Faithfulness).
-These results demonstrate that AuditEval successfully fulfills its goal of bridging the evaluation gap in audio editing by providing a scalable and repeatable MOS-style assessment tool to support model comparison and benchmarking.
-At the utterance level, the mean squared error (MSE) between AuditEval predictions and human ratings remains below 0.3 (on a 1–5 scale) across all dimensions, with the linear correlation coefficient (LCC) exceeding 0.3 in each case, indicating positive correlation (see Figure 3).
-Compared to system-level evaluation, utterance-level prediction is inherently more challenging due to increased variability.
-Nevertheless, AuditEval achieves relatively low MSE values and maintains a reasonable correlation with human judgments, demonstrating its ability to approximate human perception at finer granularity.
-Importantly, the closeness of AuditEval’s predictions to human ratings in absolute terms lays a solid foundation for downstream applications such as automatic utterance-level filtering or selection.
-Among the three dimensions, Faithfulness consistently achieves the highest performance, both at the system and utterance levels.
-This indicates that the model is relatively effective in evaluating whether the unedited content is preserved after editing.
-In contrast, Relevance and Quality are more challenging to predict accurately, as they require a deeper joint understanding of both the original and edited audio-text pairs.
-Unlike standard audio generation evaluation—which typically involves assessing a single audio clip against its corresponding text—audio editing evaluation introduces an additional layer of semantic and acoustic reasoning across paired inputs.
-To further assess the effectiveness of our proposed filtering strategy for the Pseudo-parallel Audio Editing Dataset, we compare the quality of three randomly sampled subsets (Random1-3) against a curated subset (Filtered) using standard objective metrics.
-As shown in Table 3, the Filtered subset outperforms three random baselines across most of the metrics, achieving the highest CLAP-based semantic alignment (CLAP_t: 49.43%, CLAP_a: 82.76%), as well as superior IS (12.61) and KL (0.6783) scores.
-(a) Quality
-(b) Relevance
-(c) Faithfulness
-Figure 4: Utterance-level fit between AuditEval predictions and human ratings cross three evaluation aspects.
-These results demonstrate that the proposed filtering process significantly improves semantic consistency and editing quality in the pseudo-parallel data.
-Nevertheless, certain objective quality metrics, such as FAD and FD, do not fully capture the improvements brought by our filtering strategy.
-Interestingly, the Filtered subset exhibits noticeably worse FAD and FD scores compared to the random subsets.
-This discrepancy underscores a fundamental divergence in the focus of subjective and objective evaluations: while human judgments are more attuned to high-level semantic attributes—such as relevance, coherence, and editing accuracy—objective metrics like FAD and FD measure more low-level acoustic properties and may even surpass human perception in such fine-grained distinctions.
-This distinction explains why semantically grounded metrics (e.g., CLAP, IS) align more closely with human evaluations, whereas acoustic metrics tend to diverge.
-Therefore, subjective evaluation remains a vital and complementary tool for assessing editing quality, particularly in semantically rich tasks.
-Conclusion
-In this paper, we address the longstanding challenge of evaluating audio editing systems by introducing AuditEval, a scalable framework for automatic MOS-style assessment.
-Through the construction of the AuditScore benchmark and the development of the AuditEval model, we enable accurate and efficient quality estimation across multiple dimensions.
-Furthermore, we demonstrate the practical utility of AuditEval by applying it to curate a high-quality pseudo-parallel dataset, which effectively leads to tangible improvements in data quality.
-Overall, our work lays a solid foundation for future research in both evaluation and data construction for audio editing.
-Appendix A Appendix
-Appendix B A. Details of Scoring Criteria Specification
+To address these issues, we apply AuditEval to automatically evaluate and filter synthesized editing pairs based on scoring criteria informed by AuditScore.
+This filtering is expected to improve data quality and semantic plausibility, advancing toward a practical and scalable supervised audio editing dataset (detailed filtering configurations provided in Appendix C).
+5 Experiments
+Sub-set Filtering Strategy Rule-Based Objective Metrics Human-Aligned Automatic Scores
+[…]
+Random1 48.70 12.23 0.2859 3.207 0.6913 62.72 5.770 3.601 3.664 4.965
+Random2 48.47 12.23 0.2816 3.197 0.6892 62.67 5.769 3.598 3.664 4.964
+Random3 48.76 12.20 0.2830 3.205 0.6867 62.74 5.769 3.601 3.667 4.965
+Filtered (AuditEval-ssl) 49.43 12.61 0.4587 4.356 0.6783 64.51 5.835 3.619 3.775 5.105
+Filtered (AuditEval-llm) 50.96 11.86 0.6003 8.596 0.5289 64.98 5.802 3.771 3.815 5.051
+Table 3: Comparison of pseudo-parallel data quality under different filtering strategies across objective metrics and human-aligned scores.
+5.1 Experiment Setup
+For MOS prediction, we train the proposed AuditEval family of models on the curated AuditScore dataset using a fixed 9:1 train–test split, with detailed training configurations provided in Appendix H.
+To evaluate consistency with human judgments, we report Mean Squared Error (MSE), Linear Correlation Coefficient (LCC), Spearman’s Rank Correlation Coefficient (SRCC), and Kendall’s Tau Correlation Coefficient (KTAU) at both the utterance and system levels, measuring prediction accuracy as well as rank-based agreement between predicted and ground-truth scores.
+For the task of selecting high-quality pseudo-parallel audio editing data, we further conduct exploratory experiments combining objective and automatic subjective evaluations to compare editing quality before and after model-based filtering.
+In this setting, we retain the top 15% of 59400 samples (over 3,300 hours) according to the predicted scores for comparison.
+This ratio is chosen as the minimal proportion at which the three random subsets exhibit comparable quality.
+Regarding evaluation metrics, following practice in TTA, we adopt IS, FAD, FD, and KL divergence to assess similarity between audios Liu et al.
+(2023) before and after editing.
+For objective cross-modal evaluation, we use CLAP Score to measure alignment between edited audio and the corresponding target descriptions.
+For automatic subjective evaluation, we employ Human-CLAP Takano et al.
+(2025), which aligns CLAP-based scores with human preferences, as well as AudioBox to provide human-aligned aesthetic quality assessments of edited audio from four perceptual dimensions: Production Quality (PQ), Content Enjoyment (CE), Content Usefulness (CU), and Production Complexity (PC).
+5.2 Results and Analysis
+Effectiveness of AuditEval.
+As shown in Table 2 and the right columns of Fig.
+1, both AuditEval-ssl and AuditEval-llm achieve practical and competitive performance in predicting audio editing quality.
+At the system level, which is commonly used to assess overall model performance and rank different systems, AuditEval-ssl attains a high LCC of around 0.6 on the Quality dimension, together with a low MSE below 0.1 on the 1–5 rating scale.
+Meanwhile, AuditEval-llm achieves remarkably high correlations of 0.76 and 0.91 on the Relevance and Faithfulness dimensions, respectively, while maintaining MSE values below 0.1, indicating substantially stronger alignment with human perception and clearly outperforming conventional objective metrics.
+Compared with system-level evaluation, utterance-level prediction is inherently more challenging due to increased variability and noise at the individual sample level.
+Nevertheless, across all dimensions, both AuditEval-ssl and AuditEval-llm maintain relatively low MSE values (approximately 0.2 on the 1–5 scale) and preserve reasonable positive correlations (ranging from 0.38 to 0.72) with human judgments.
+In particular, AuditEval-llm achieves correlations of 0.47 for Relevance and 0.72 for Faithfulness, excelling in semantic related dimensions.
+These results lay a solid foundation at a finer granularity for subsequent downstream utterance-level data filtering and selection.
+AuditEval-ssl vs. AuditEval-llm.
+A comparison between AuditEval-ssl and AuditEval-llm (Table 2 and Fig.
+1 right) shows that AuditEval-llm achieves state-of-the-art performance across most dimensions and metrics.
+In particular, on the Relevance and Faithfulness dimensions, which are closely related to semantic understanding, AuditEval-llm demonstrates notably stronger judgment capability.
+This advantage is likely attributable to the reasoning abilities and intelligence of large audio language models.
+From an acoustics-oriented perspective, AuditEval-ssl exhibits comparable, and in some metrics even superior, prediction performance on Quality dimension.
+This aligns with the general observation that audio-specific self-supervised learning models are particularly effective at extracting fine-grained acoustic representations, especially high-frequency cues that influence audio quality.
+Effectiveness of Filtering.
+As shown in Table 3, the filtered data, selected using either AuditEval-ssl or AuditEval-llm, consistently exhibits improved performance across all human-aligned scores, as well as on CLAP_Score, which is currently one of the most reliable objective metrics for audio editing.
+Notably, the subjective evaluation models used in our experiments, namely Human-CLAP and AudioBox, are entirely independent of AuditEval in terms of both data and model design.
+This fully in-the-wild evaluation setting provides strong evidence that the proposed filtering strategy leads to consistent and reliable quality improvements.
+Objective Metrics vs. Human-aligned Scores.
+As shown in Table 3, improvements in human-aligned scores are not consistently observed in objective metrics, contrary to expectations.
+Instead, metrics such as FAD and FD, which are commonly used to measure similarity between original and edited audios (faithfulness), exhibit noticeable degradation after filtering.
+This discrepancy reflects a fundamental difference between subjective and objective evaluation criteria.
+Human judgments tend to emphasize overall perceptual quality, including coherence, enjoyment, usability, and the semantic correctness of editing operations.
+In contrast, objective metrics such as FAD and FD primarily focus on low-level acoustic signal variations that are often perceptually insignificant to human listeners.
+These observations highlight the limitations of objective metrics and underscore the advantages of human-aligned evaluation, particularly for the complex and semantically rich audio editing scenarios.
+6 Conclusion
+In this paper, we address the long-standing challenges of data scarcity and unreliable evaluation in audio editing by grounding both evaluation and data construction in human judgments.
+We systematic construct AuditScore, the first human-annotated dataset for MOS-style audio editing assessment, then develop AuditEval-ssl and AuditEval-llm to study two different paradigms for automatic subjective evaluation and high-quality dataset mining, providing both practical tools and methodological insights.
+Through comprehensive experiments, we reveal key challenges in audio editing evaluation and data curation, and highlight the role of human-aligned evaluation in improving evaluation and data quality.
+Appendix A Details of Scoring Criteria Specification
 Here, we present details about Scoring Criteria Specification.
@@
 Quality score reflects how well the edited audio aligns with the editing instruction and target description, in terms of semantic correctness and execution:
-$\bullet$ 5 - About the Same/Even Better: The edited audio is on par with or even better than the original in terms of intelligibility, fluency, and background noise.
-$\bullet$ 4 - Slightly Worse: At most one aspect (e.g., mild mechanical artifact or minor increase in noise) is slightly degraded, but overall listening quality remains unaffected.
-$\bullet$ 3 - Noticeably Worse: At least two aspects show moderate degradation, or one aspect is clearly degraded.
+  • 5- About the Same/Even Better: The edited audio is on par with or even better than the original in terms of intelligibility, fluency, and background noise.
+  • 4- Slightly Worse: At most one aspect (e.g., mild mechanical artifact or minor increase in noise) is slightly degraded, but listening quality remains unaffected.
+  • 3- Noticeably Worse: At least two aspects show moderate degradation, or one aspect is clearly degraded.
 Artifacts such as robotic tone, discontinuities, or increased noise may affect understanding in some segments but remain tolerable.
-$\bullet$ 2 - Much Worse: Significant degradation in at least two aspects, with noticeable artifacts severely affecting naturalness and intelligibility.
-$\bullet$ 1 - Badly Damaged: Severe issues across all aspects, such as overwhelming noise, harsh distortions, or unintelligible content, rendering the audio difficult to comprehend.
+  • 2- Much Worse: Significant degradation in at least two aspects, with noticeable artifacts severely affecting naturalness and intelligibility.
+  • 1- Badly Damaged: Severe issues across all aspects, such as overwhelming noise, harsh distortions, or unintelligible content, rendering the audio difficult to comprehend.
 Relevance score reflects how well the edited audio aligns with the editing instruction and target description, in terms of semantic correctness and execution:
-$\bullet$ 5 - Excellent Match: The edit perfectly matches the described intent.
+  • 5- Excellent Match: The edit perfectly matches the described intent.
 For Addition, the added event feels natural and contextually appropriate.
@@
 For Replacement, the substituted content is fluent and coherent, maintaining the overall auditory experience.
-$\bullet$ 4 - Good Match: Overall alignment is correct, but there may be minor imperfections, such as slightly abrupt additions or faint remnants after deletion.
-$\bullet$ 3 - Partial Match: The general category or type of edit is correct, but details may be awkward or somewhat misaligned with the description.
-$\bullet$ 2 - Poor Match: The edit loosely corresponds to the intent, but contains semantic inaccuracies or unnatural execution that hampers comprehension.
-$\bullet$ 1 - Mismatch: The edited content fails to reflect the described intent, with missing, incorrect, or misleading modifications.
+  • 4- Good Match: Overall alignment is correct, but there may be minor imperfections, such as slightly abrupt additions or faint remnants after deletion.
+  • 3- Partial Match: The general category or type of edit is correct, but details may be awkward or somewhat misaligned with the description.
+  • 2- Poor Match: The edit loosely corresponds to the intent, but contains semantic inaccuracies or unnatural execution that hampers comprehension.
+  • 1- Mismatch: The edited content fails to reflect the described intent, with missing, incorrect, or misleading modifications.
 Faithfulness criterion evaluates how well the unedited portions of the audio are preserved in terms of rhythm, sequence, loudness, and overall style:
-$\bullet$ 5 - Perfectly Faithful: Unedited segments are completely preserved. Prosody, order, and audio style remain fully consistent with the original.
-$\bullet$ 4 - Mostly Faithful: Slight variation in loudness or subtle stylistic shifts may occur, but prosody and content are largely preserved.
-$\bullet$ 3 - Partially Faithful: Noticeable changes in rhythm or sequence are present, though the overall style remains similar. Minor alterations in the unedited content may be perceptible.
-$\bullet$ 2 - Low Faithfulness: Major changes in prosody, ordering, or style, with only a few elements traceable to the original.
-$\bullet$ 1 - Not Faithful: The edited audio bears little or no resemblance to the original; unedited content is altered significantly or missing altogether.
-Appendix C B. Details of Pseudo-Parallel Audio Editing Dataset Construction.
+  • 5- Perfectly Faithful: Unedited segments are completely preserved. Prosody, order, and audio style remain fully consistent with the original.
+  • 4- Mostly Faithful: Slight variation in loudness or subtle stylistic shifts may occur, but prosody and content are largely preserved.
+  • 3- Partially Faithful: Noticeable changes in rhythm or sequence are present, though the overall style remains similar. Minor alterations in the unedited content may be perceptible.
+  • 2- Low Faithfulness: Major changes in prosody, ordering, or style, with only a few elements traceable to the original.
+  • 1- Not Faithful: The edited audio bears little or no resemblance to the original; unedited content is altered significantly or missing altogether.
+Appendix B Details of Pseudo-Parallel Audio Editing Dataset Construction.
 Inspired by AUDIT and InstructME, we also adopt an audio mixing strategy to construct our Pseudo-Parallel Audio Editing Dataset.
 The goal of the dataset construction process is to generate a set of audio pairs comprising (A, B, C, A + B, A + C), which serves as the foundation for subsequent audio editing tasks.
-For audio types A, we collected over 45,000 audio prompts from the AudioCaps dataset (Kim et al.
-2019).
+For audio types A, we collected over 45,000 audio prompts from the AudioCaps dataset Kim et al.
+[2019].
 For audio types B and C, we select prompts corresponding to common single-audio events, covering eight broad categories—NaturalSounds, Transportation, IndoorActivities, OutdoorActivities, Animals, Music, Speech and HumanSounds—spanning over 100 distinct sound types.
@@
 Then, for each audio prompt pair (B, C), we randomly select 20 audio prompts from the 45,000+ A audio prompts to form combinations, denoted as (A, B, C).
-We utilize TangoFlux(Hung et al. 2024) to generate audio based on the (A, B, C) prompts, using a step of 50 and a guide scale of 4.5. The generation rules are as follows:
+We utilize TangoFlux Hung et al. [2024] to generate audio based on the (A, B, C) prompts, using a step of 50 and a guide scale of 4.5. The generation rules are as follows:
   • Audio A: The generated audio has a fixed length of 10 seconds.
@@
 Specifically, (Add, A, A + B) and (Add, A, A + C) constitute two training examples for the add task; (Delete, A + B, A) and (Delete, A + C, A) constitute two training examples for the delete task; (Replace, A + B, A + C) and (Replace, A + C, A + B) constitute two training examples for the replace task.
-Appendix D C. Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
-Table 5: Number of Samples After Each Round of Selection
+Figure 3: Score prediction distributions on the large-scale Pseudo-Parallel Audio Editing Dataset, with addition, deletion, and modification operations shown from top to bottom.
+Figure 4: Score prediction distributions on the large-scale Pseudo-Parallel Audio Editing Dataset, with addition, deletion, and modification operations shown from top to bottom.
+Appendix C Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
+Table 5: Number of Samples After Each Round of Selection(ssl)
 Selection Stage Add Delete Replace
@@
 After Round 3 30000 30000 30000
+Table 6: Number of Samples After Each Round of Selection(llm)
+Selection Stage Add Delete Replace
+Original number 198000 198000 198000
+After Round 1 99000 99000 99000
+After Round 2 68551 76358 68522
+After Round 3 30000 30000 30000
 Here, we present the filtering strategy of constructing High-Quality Pseudo-Parallel Audio Editing Dataset.
-In AppendixC, the construction of the Pseudo-Parallel Audio Editing Dataset is presented, comprising 60,000 training instances for each of the three editing tasks: Addition, Deletion, and Replacement.
-We employ our proposed model, AuditEval, to evaluate the dataset across three dimensions(Quality, Relevance, and Faithfulness).
+In AppendixC, the construction of the Pseudo-Parallel Audio Editing Dataset is presented, comprising 30,000 training instances for each of the three editing tasks: Addition, Deletion, and Replacement.
+We employ our proposed model, AuditEval, to evaluate the dataset across three dimensions, with the score distributions illustrated in Fig.
+3 and Fig.
+4.
 Subsequently, based on the scores of these three dimensions, we perform three rounds of filtering on the dataset.
-The filtering criteria are detailed as follows:
-  • Stage 1: Faithfulness-based filtering. All samples are sorted by their predicted faithfulness scores in descending order. We retain the top 30% of samples as determined by this ranking.
-  • Stage 2: Quality thresholding. Among the retained samples, we further filter out those with predicted quality scores below 3.4.
+The filtering criteria are as follows:
+  • Stage 1: Faithfulness-based filtering. All samples are sorted by their predicted faithfulness scores in descending order.
+  • Stage 2: Quality thresholding. Among the retained samples, we further filter out those with predicted quality scores boardline.
   • Stage 3: Relevance-based selection. The remaining samples are ranked by predicted relevance scores, and the top 30,000 samples are selected to form the final dataset.
-In practice, we use a quality threshold of 3.40, a faithfulness threshold of 3.2156/3.2813/3.2662 and a relevance threshold of 3.1592/3.4844/3.3441.
-Table5 illustrates the selection process, where the chosen thresholds approximately eliminate half of the data at each round.
-Appendix E D. User Interface for Subjective Evaluation
+For AuditEval-ssl, we adopt a quality threshold of 3.40, faithfulness thresholds of 3.216 / 3.281 / 3.266, and relevance thresholds of 3.159 / 3.484 / 3.344 for the three successive filtering rounds.
+For AuditEval-llm, the corresponding thresholds are set to 3.70 for quality, 3.578 / 3.641 / 3.453 for faithfulness, and 3.500 / 3.609 / 3.469 for relevance.
+Tables 5 and 6 illustrate the multi-round selection process for the two evaluators, respectively.
+In each round, the selected thresholds are designed to remove approximately half of the remaining samples, progressively retaining higher-quality pseudo-parallel data.
+Appendix D User Interface for Subjective Evaluation
+Figure 5: User Interface for Subjective Evaluation.
 To ensure consistency and ease of use during the manual scoring process, we designed a custom web-based interface for subjective evaluation.
 As shown in Figure 5, the interface presents the original audio, the edited audio, the original prompt, the edited prompt, and, if available, the edit instruction.
-Annotators are instructed to score each sample along three dimensions—Quality, Relevance, and Faithfulness—following the same evaluation criteria defined in the AuditScore dataset.
+Annotators are instructed to score each sample along three dimensions: Quality, Relevance, and Faithfulness, following the same evaluation criteria defined in the AuditScore dataset.
 This standardized interface streamlines the annotation workflow and helps reduce variability in human ratings.
-Figure 5: User Interface for Subjective Evaluation.
-Appendix F E. Computational Experiments Details
-All experiments were conducted on a single consumer-grade GPU: NVIDIA GeForce RTX 4090. To ensure reproducibility, we fixed all sources of randomness with a global random seed set to 1984.
-The AuditEval model was trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$.
-A learning rate scheduler ReduceLROnPlateau was employed with the following parameters: mode=’min’, factor=0.8, patience=5, and verbose=True.
-The training was performed with a batch size of 256 for training and 64 for validation.
-The model consists of three independent regression heads corresponding to the Quality, Relevance, and Faithfulness dimensions, trained for different numbers of epochs: Quality: trained for 50 epochs; Relevance: trained for 30 epochs; Faithfulness: trained for 35 epochs.
-Appendix G F. Licenses for assets
-This work uses publicly available resources.
-AudioCaps and the audio editing inversion code are released under the MIT License.
-CLAP pretrained weights are under the CC0-1.0 License.
-All assets developed in this work—including datasets, checkpoints, training protocols, and code—will be released under the CC BY 4.0 License.
-Appendix H G. Ethics Statement
+Appendix E Computational Experiments Details
+All experiments were conducted on a single NVIDIA H800 GPU. To ensure reproducibility, we fixed all sources of randomness with a global random seed set to 1984.
+The AuditEval-ssl model was trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$.
+We employed a ReduceLROnPlateau scheduler with mode=’min’, factor=0.8, and patience=5.
+Training was conducted with a batch size of 256 for training and 64 for validation.
+The model comprises three independent regression heads for Quality, Relevance, and Faithfulness, which were trained for 50, 30, and 35 epochs, respectively.
+AuditEval-llm was trained within the ms-swift framework as a regression-based sequence classification model that predicts three continuous scores corresponding to Quality, Relevance, and Faithfulness (num_labels=3, problem_type=regression), with chat templates disabled to ensure a unified input format.
+We adopted LoRA-based parameter-efficient fine-tuning, applying LoRA to all linear layers with a rank of 8 and a scaling factor (alpha) of 32, while freezing the ViT backbone.
+The model was trained for 6 epochs using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a warmup ratio of 0.05, under mixed-precision training with bfloat16.
+The per-device batch size was set to 8 for training and 2 for evaluation, with gradient accumulation over 8 steps and a maximum input sequence length of 4096 tokens.
+Appendix F Licenses for assets
+All assets developed in this work, including datasets, checkpoints, training protocols, and code, will be released under the CC BY 4.0 License.
+Appendix G Ethics Statement
 We employed professional annotators via a third-party commercial provider to ensure consistent and reliable subjective evaluations.
 Annotators were fairly compensated above local minimum wage in accordance with regional labor laws.
-They received clear instructions on task objectives, scoring criteria (Quality, Relevance, Faithfulness), and example cases prior to annotation.
+They received clear instructions on task objectives, scoring criteria, and example cases prior to annotation.
 No personal or sensitive information was collected, and no individual-level data beyond task performance was recorded.
-The provider ensured fair labor practices, contractual compliance, and data confidentiality.
-As annotators were not research subjects but task workers evaluating system outputs, IRB approval was not required.
-Appendix I H. Evaluation scores Across Editing Operations (Add / Delete / Replace)
+LLMs like GPT-5 were used appropriately in this work, primarily for language polishing.
+We ensured that all content was human-authored and carefully reviewed, with no reliance on LLMs for generating scientific claims or experimental results, thereby avoiding potential AI hallucinations.
+Appendix H Details of the AuditEval-llm Prompt
+Figure 6 and Figure 7 illustrate the prompts used in AuditEval-llm, which instruct the model to act as a professional audio editing quality evaluator.
+Prompt You are a professional audio editing quality evaluator.
+You will be given: (1) the original audio, (2) the edited audio, (3) a textual description of the original audio (original_caption), (4) a textual description of the target edited audio (target_caption).
+The semantic difference between original_caption and target_caption implicitly defines the intended audio edit.
+Your task is to evaluate how well the edited audio reflects this semantic change, while maintaining high audio quality and preserving unedited content.
+Evaluate the edited audio from three perspectives: Quality, Relevance, and Faithfulness.
+Each score should be an integer from 1 to 5 (higher is better).
+Output only three integer scores, without explanations.
+1.
+Quality (1-5): Evaluate the perceptual quality of the edited audio compared to the original, including intelligibility, fluency, naturalness, and background noise.
+- 5: As good as or better than the original, with clear, fluent, and natural audio.
+- 4: Slight degradation in at most one aspect, but overall quality remains acceptable.
+- 3: Noticeable degradation in multiple aspects or clear degradation in one aspect, but still understandable.
+Figure 6: Examples AuditEval-llm Promps(I)
+Prompt - 2: Significant artifacts that seriously harm naturalness or intelligibility.
+- 1: Severely damaged or unintelligible audio.
+2.
+Relevance (1-5): Evaluate how well the edited audio matches the semantic content described in target_caption, relative to original_caption.
+Focus only on the parts that are different between original_caption and target_caption.
+- 5: The edited audio accurately and naturally reflects the target_caption, with correct and complete semantic changes.
+- 4: The main semantic change is correctly reflected, with only minor imperfections.
+- 3: The general semantic intent is correct, but details are partially misaligned or awkward.
+- 2: The semantic change is loosely reflected, with clear errors or unnatural execution.
+- 1: The edited audio does not match the target_caption.
+3.
+Faithfulness (1-5): Evaluate how well the audio content that should remain unchanged (according to the captions) is preserved in the edited audio, in terms of rhythm, order, loudness, prosody, and style.
+Do not penalize the semantic changes required by target_caption.
+- 5: Unchanged content is perfectly preserved and indistinguishable from the original.
+- 4: Mostly preserved, with only minor variation.
+- 3: Noticeable changes in rhythm or prosody, but overall style remains similar.
+- 2: Major changes in unedited content with limited resemblance.
+- 1: Unedited content is heavily altered or missing.
+Original Caption: {original_caption} Target Caption: {target_caption}
+Figure 7: Examples AuditEval-llm Promps(II)
+Appendix I Evaluation scores Across Editing Operations (Add / Delete / Replace)
+This section presents a breakdown of evaluation score across different audio editing operations, including Add, Delete, and Replace, as shown in following table:
 Algorithm TTA backbone Timesteps System ID Operation Quality score Relevance score Faithfulness score
@@
 Rep 3.840 3.392 3.450
-ZETA
-Tango2 75 sys9 Add 3.798 3.530 3.842
+ZETA Tango2 75 sys9 Add 3.798 3.530 3.842
 Del 3.620 3.154 3.604